### v1

# BConformeR: A Conformer Based on Mutual Sampling for Unified Prediction of Continuous and Discontinuous Antibody Binding Sites

Abstract
Accurate prediction of antibody-binding sites (epitopes) on antigens is crucial for vaccine design, immunodiagnostics, therapeutic antibody development, antibody engineering, research into autoimmune and allergic diseases, and for advancing our understanding of immune responses. Despite in silico methods that have been proposed to predict both linear (continuous) and conformational (discontinuous) epitopes, they consistently underperform in predicting conformational epitopes. In this work, we propose a conformer-based model trained on antigen sequences derived from 1,080 antigen-antibody complexes, leveraging convolutional neural networks (CNNs) to extract local features and Transformers to capture long-range dependencies within antigen sequences. Ablation studies demonstrate that CNN enhances the prediction of linear epitopes, and the Transformer module improves the prediction of conformational epitopes. Experimental results show that our model outperforms existing baselines in terms of PCC, ROC-AUC, PR-AUC, and F1 scores on conformational epitopes.
Keywords B-cell Epitope Prediction; Conformer; Antigen–Antibody Interface
1 Introduction
B cells are key effectors of the adaptive immune system, coordinating responses to pathogenic and aberrant signals. Each B cell expresses a single clonotype of B-cell receptor (BCR), a membrane-bound antibody that interacts with specific antigens. The binding sites on the antigen—known as the epitope—can be linear (continuous) or conformational (discontinuous), and the corresponding binding sites on the antibody are called paratopes. Linear epitopes are formed by linear stretches of amino acid residues, while conformational epitopes are formed by distant residues brought together in 3D through folding.
Figure 1: Overview of epitope classification and BConformerR prediction. (A) Epitopes, i.e., Antibody-binding sites, include conformational fragments (n=2,632) and linear regions (n=2,388). Data derived from 1,432 antigen records. (B) Epitope prediction workflow: Antigen inputs are processed to predict linear and conformational epitopes, which enable applications in vaccine design and diagnostic development.
Identifying epitopes is critical for antibody engineering koz:18 ; ji:22 , disease diagnosis va:22 ; ta:24 , and vaccine optimization dv:24 ; ba:25 . Experimental techniques such as X-ray crystallography and cryo-electron microscopy provide high-resolution antigen-antibody (Ag–Ab) interactions but are resource-intensive crys:21 ; cryoEM:25 . Phage display is faster but lacks atomic-level precision pd:22 . Therefore, many in silico methods have been developed to predict epitopes.
Sequence-based methods such as SEMA-1D SEMA ; SEMA2.0 , BepiPred BepiPred1.0 ; BepiPred2.0 ; BepiPred3.0 and SEPPA SEPPA3.0 are widely used due to the availability of antigen sequences, with SEMA-1D 2.0 SEMA2.0 and BepiPred-3.0 BepiPred3.0 notably improving performance by incorporating representations from the protein language model ESM-2 ESM-2pre ; ESM-2 . Structure-based methods, such as SEMA-3D SEMA ; SEMA2.0 , DiscoTope DiscoTope2.0 ; DiscoTope3.0 , ScanNet ScanNet and Epitope3D Epitope3D , take the residue geometry and surface accessibility into consideration, but remain constrained by the limited availability of high-quality antigen structures. Among these methods, DiscoTope-3.0 DiscoTope3.0 achieves the best performance in the benchmark evaluation.
Figure 2: BConformeR overview. (A) Preparation pipeline of the data. 1300 antigen-antibody FASTA files were processed using AlphaFold Multimer v3 to predict structures, with top-ranked structure for each was selected to extract epitopes (right); Antigen sequences were embedded by ESM-2, producing embeddings of dimension 1024×1280 (left). (B) Conformer architecture. Stem module applies convolution and max-pooling to the original ESM-2 embeddings; The Conv block and Trans block are based on ResNet bottleneck and Vision Transformer modules respectively; Feature coupling units (FCU-Up and FCU-Down) connect the Conv and Trans blocks, with 11 FCU layers in total; The classifier linearly fuses features from both Conv and Trans branches.
However, epitope prediction remains a challenging task, with existing methods underperform on conformational ones. Moreover, there is a lack of hybrid backbone models to effectively handle both linear and conformational epitopes. A key source of this complexity stems from the nature of Ag-Ab interfaces, where linear epitopes are clustered along the sequence, and conformational epitopes are often spatially proximate but dispersed across different regions of the linear sequence OverviewEp . To address this multi-scale complexity, we designed the Conformer architecture, which integrates convolution layers to learn local residue-level features and self-attention mechanisms to capture long-range dependencies within antigen sequences Conformer20 ; Conformer21 .
Here, we propose BConformeR, a Conformer-based model for linear and conformational BCR epitope prediction. Antigen sequences from 1,180 Ag-Ab complexes were embedded using ESM-2, with 1,080 used for training and 100 for evaluation. In a blind test of 24 antigen records, BConformeR outperforms existing baselines—SEPPA-3.0, BepiPred-3.0, SEMA-1D 2.0 and DiscoTope-3.0—across multiple metrics, including PCC, ROC-AUC, PR-AUC, and F1 score on conformational epitopes. To further dissect the model’s design, we ablated ResNet bottlenecks and ViT modules from Conformer, showing that convolutional layers improve linear epitope prediction, and Transformer modules are critical for capturing discontinuous epitope patterns.
2 Related Work
2.1 Antibody-agnostic prediction models
Numerous models, trained on antigen–epitope pairs, have been proposed and refined to enhance epitope prediction. SEPPA-3.0 was the epitope prediction tool specifically designed for glycoproteins, based on the logistic regression classifier SEPPA3.0 . BepiPred-3.0 used embeddings from pretrained protein language models ESM-2 as input to a feedforward neural network (FFNN) to predict both linear and conformational B-cell epitopes BepiPred3.0 . Epitope3D modeled antigen surfaces as residue-level graphs and extracted distance-based graph signatures to train classifiers for conformational B-cell epitope prediction, with AdaBoost selected as the backbone Epitope3D ; AdaBoost . ScanNet was an end-to-end geometric deep learning model that learned interpretable residue-level features directly from 3D structural neighborhoods ScanNet . SEMA-1D 2.0 was constructed by adding a fully-connected linear layer on an ensemble of five ESM-2 models, while SEMA-3D 2.0 was developed using an ensemble of five pre-trained Structure-aware Protein language models (SaProt) SEMA2.0 ; SaProt . DiscoTope-3.0 utilized inverse folding representations from ESM-IF1 and was trained on both predicted and solved structures using a positive-unlabelled ensemble strategy, enabling structure-based B-cell epitope prediction at scale DiscoTope3.0 ; ESM-IF1 .
However, existing models rarely incorporate both local features and long-range dependencies within antigen sequences, showing limited performance on conformational epitopes. Therefore, we adopted a Conformer based on mutual sampling, employing CNNs to extract local features and Transformers to capture long-range dependencies within antigen sequences.
2.2 Antibody-aware evaluation models
Despite advances in epitope prediction, evaluating Ag-Ab binding confidence remains challenging, which is crucial for vaccine selection. AbEpitope-1.0, a recent method for predicting antibody targets by use of AlphaFold and inverse folding, significantly improving the prediction accuracy of Ag-Ab interfaces. They constructed negative Ag-Ab samples (false binding pairs) by swapping antibody chains across real complexes in the dataset. Each sample was labeled as either true or false to train the model to distinguish authentic Ag-Ab interactions from decoys. The final model, AbEpiTarget-1.0, achieved superior performance in antibody-sensitive scoring for target prediction AbEpitope .
In this work, we utilize the real Ag-Ab complex dataset provided by AbEpitope-1.0, employing its epitope identification pipeline to ensure the data quality (Fig. 2A).
3 Method
3.1 Epitope Extraction
3.1.1 Epitope identification
During immune recognition, epitopes are typically composed of multiple spatially adjacent residues. To capture this interaction, the notions of residue-neighbor and residue-patch were introduced. A residue-neighbor is defined when the minimum distance between any atoms of two residues is less than 4 Å, and a residue-patch refers to a group of residues whose atoms lie within 10 Å of a central residue epitope_id . In our study, we focused on the residue-neighbor relationship to capture the tight spatial interactions, which is given as:
$\min_{a\in r_{i},\,b\in r_{j}}\|a-b\|<4\,\text{\AA }$, (1)
where $r_{i}$ and $r_{j}$ are residues from the antigen and antibody chains respectively, and $a$, $b$ represent the heavy atoms within these residues.
3.1.2 Epitope classification
Generally, epitopes are separated into multiple Regions (R) and Fragments (F) for each antigen sequence. Regions are defined as linear stretches of antigen sequence having at least three antibody-contacting residues. Gaps between contacting residues are allowed, and a gap size of up to three non-contacting residues is chosen on the basis of the structure of $\alpha$-helices, which enables residues on the same helical face to be included epitope_cl . Fragments, which do not contain enough antibody-contacting residues to qualify as a region, are considered discontinuous in our study (Fig. 1A).
For real and predicted annotations, linear and discontinuous epitopes are extracted from regions and fragments, respectively. Ambiguous cases occur when the real and predicted labels differ, especially when a residue is annotated as part of a region in the ground truth but as part of a fragment in the prediction, or vice versa. To address this, all classifications are determined based on the ground-truth labels. That is, for a real label $r\in\{N,L,D\}$ and predicted label $p\in\{N,L,D\}$, the final classification $c(r,p)$ is defined as:
$c(r,p)=\begin{cases}L,&(r,p)\in\{(L,\cdot),(N,L)\}\\ D,&(r,p)\in\{(D,\cdot),(N,D)\}\\ N,&(r,p)=(N,N)\end{cases}$ (2)
where $N$, $L$, and $D$ denote non-epitope, linear epitope and discontinuous epitope, respectively. Performance metrics (e.g., F1) are then calculated for each type of epitopes.
stage output CNN Branch FCU Transformer Branch
stem 1024 7, 320
3 max pooling
s1 1024 $\left[\begin{array}[]{l}\text{1, 160}\\[-1.00006pt] \text{3, 160}\\[-1.00006pt] \text{1, 640}\end{array}\right]$ - 5, 1536 $\times$1
$\left[\begin{array}[]{c}\text{\text{MHSA-12}, 1536}\\[-1.00006pt] \text{1, 3072}\\[-1.00006pt] \text{1, 1536}\end{array}\right]$
$\left[\begin{array}[]{l}\text{1, 160}\\[-1.00006pt] \text{3, 160}\\[-1.00006pt] \text{1, 640}\end{array}\right]$ $\times$3
$1,1536]\longrightarrow$
$\left[\begin{array}[]{c}\text{\text{MHSA-12}, 1536}\\[-1.00006pt] \text{1, 3072}\\[-1.00006pt] \text{1, 1536}\end{array}\right]$
$\left[\begin{array}[]{l}\text{1, 160}\\[-1.00006pt] \text{3, 160}\\[-1.00006pt] \text{1, 640}\end{array}\right]$
$\longleftarrow[1,160]$
s2 1024 $\left[\begin{array}[]{l}\text{1, 320}\\[-1.00006pt] \text{3, 320}\\[-1.00006pt] \text{1, 1280}\end{array}\right]$ $\times$4
$1,1536]\longrightarrow$
$\left[\begin{array}[]{c}\text{\text{MHSA-12}, 1536}\\[-1.00006pt] \text{1, 3072}\\[-1.00006pt] \text{1, 1536}\end{array}\right]$
$\left[\begin{array}[]{l}\text{1, 320}\\[-1.00006pt] \text{3, 320}\\[-1.00006pt] \text{1, 1280}\end{array}\right]$
$\longleftarrow[1,320]$
s3 1024 $\left[\begin{array}[]{l}\text{1, 640}\\[-1.00006pt] \text{3, 640}\\[-1.00006pt] \text{1, 2560}\end{array}\right]$ $\times$4
$1,1536]\longrightarrow$
$\left[\begin{array}[]{c}\text{\text{MHSA-12}, 1536}\\[-1.00006pt] \text{1, 3072}\\[-1.00006pt] \text{1, 1536}\end{array}\right]$
$\left[\begin{array}[]{l}\text{1, 640}\\[-1.00006pt] \text{3, 640}\\[-1.00006pt] \text{1, 2560}\end{array}\right]$
$\longleftarrow[1,640]$
[…]
$\alpha\times\mathrm{Conv}_{\mathrm{cls}}+(1-\alpha)\times\mathrm{Trans}_{\mathrm{cls}}$
Table 1: Architecture of BConformeR, a 1D adaptation of Conformer Conformer21 . The MHSA-12 indicates multi-head self-attention with 12 heads operating on 1536-dimensional embeddings (128 dimensions per head). $\mathrm{Conv}_{\mathrm{cls}}$ denotes the 1D convolutional classifier operating on CNN features, and $\mathrm{Trans}_{\mathrm{cls}}$ denotes the linear classifier operating on Transformer outputs. The final prediction is obtained by a weighted fusion of the two.
3.2 Conformer
3.2.1 Overview
To take advantage of both local and global representations for epitope prediction, we construct BConformeR, a dual-branch architecture adapted from Conformer Conformer21 , as shown in Table 1. BConformeR consists of a shared convolutional stem, a CNN branch, a Transformer branch, Feature Coupling Units (FCUs) for bidirectional feature exchange, and two per-position classifiers (Fig. 2B). Since our task is per-residue epitope prediction, all antigen sequences are padded to a fixed length of $L$ (e.g., 1024) and this length is preserved throughout the model.
The stem consists of a 1D convolution (kernel size 7) and a 1D max pooling (kernel size 3), designed to capture contextual features for each residue. And the embedding dimension $C$ (e.g., 1280) is reduced to $C/4$ channels after the stem. In the first block, both branches process identical initial features and therefore no feature coupling is applied. From the second block onward, FCUs are inserted to progressively fuse local and global representations. The input is processed by alternating CNN and Transformer blocks, where at each stage, CNN-derived local features are passed to the Transformer branch, and Transformer-derived global representations are fed back to the CNN branch. The CNN branch gradually expands to $2C$ channels, and the Transformer branch maintains $D$-dimensional (e.g., 1536) token embeddings at the end.
Finally, a 1D convolutional classifier and a linear classifier are applied to the outputs of the CNN and Transformer branches, respectively, producing per-residue logits $\mathbf{y}_{\text{CNN}},\mathbf{y}_{\text{Trans}}\in\mathbb{R}^{L\times 2}$. These are linearly fused and converted to per-residue class probabilities using the softmax function:
$\mathbf{p}_{i}=\mathrm{softmax}\left(\alpha\mathbf{y}_{\text{CNN},i}+(1-\alpha)\mathbf{y}_{\text{Trans},i}\right),$ (3)
where $\mathbf{y}_{i}\in\mathbb{R}^{2}$ denotes the logits for the $i$-th residue, and $\alpha$ is the weighted parameter.
3.2.2 CNN branch
The CNN branch is composed of three stages after the stem, each consisting of repeated convolutional blocks based on the ResNet bottleneck design ResNet . Given an input feature sequence $\mathbf{x}\in\mathbb{R}^{C^{\prime}\times L}$, a bottleneck performs three sequential convolutions with a residual connection applied to $\mathbf{x}$:
$\mathbf{z}_{1}$ $=\sigma\left(\text{BN}_{1}\left(\text{Conv}_{1\times 1}(\mathbf{x})\right)\right),$ (4)
$\mathbf{z}_{2}$ $=\sigma\left(\text{BN}_{2}\left(\text{Conv}_{3\times 1}(\mathbf{z}_{1})\right)\right),$
$\mathbf{z}_{3}$ $=\text{BN}_{3}\left(\text{Conv}_{1\times 1}(\mathbf{z}_{2})\right),$
$\mathbf{y}$ $=\sigma\left(\mathbf{z}_{3}+\mathcal{F}(\mathbf{x})\right),$
where $\sigma(\cdot)$ denotes the ReLU activation, BN represents 1D batch normalization, and
$\mathcal{F}(\mathbf{x})=\begin{cases}\mathbf{x}.&\text{if shape matches}\\ \text{BN}_{\text{res}}\left(\text{Conv}_{1\times 1}(\mathbf{x})\right).&\text{otherwise}\end{cases}$
The channel dimension is first downsampled to half, then expanded by a factor of 4 within each bottleneck. The first convolution block contains a single bottleneck, while subsequent blocks include two bottlenecks: one applied before the Transformer branch, and the other afterward. Since convolution kernels slide over representation sequences with overlap, the CNN blocks are well-suited to capture residue-level contextual information, providing local features for the Transformer branch.
3.2.3 Transformer branch
The Transformer branch processes residue embeddings using 12 repeated Transformer blocks. Each block follows the standard architecture introduced in ViT ViT , but adapted to 1D protein sequences. As shown in Table 1, each block contains a multi-head self-attention module and a two-layer feed-forward MLP block. Both submodules are preceded by Layer Normalization (LN) with residual connections. Specifically, given the input sequence $\mathbf{x}\in\mathbb{R}^{L\times D}$, a single Transformer block updates the representation as follows:
$\mathbf{z}_{1}$ $=\mathbf{x}+\mathrm{DropPath}\left(\mathrm{MHSA}\left(\mathrm{LN}_{1}(\mathbf{x})\right)\right),$ (5)
$\mathbf{z}_{2}$ $=\mathbf{z}_{1}+\mathrm{DropPath}\left(\mathrm{MLP}\left(\mathrm{LN}_{2}(\mathbf{z}_{1})\right)\right),$
where $\mathrm{MHSA}(\cdot)$ denotes the multi-head self-attention mechanism with $h$ (e.g., 12) heads, and $\mathrm{MLP}(\cdot)$ consists of a linear up-projection to $2D$ dimensions, a GELU activation, followed by a linear down-projection to $D$. The $\mathrm{DropPath}$ operator implements stochastic depth regularization.
Since the CNN branch already encodes positional information, no explicit positional embeddings are added. Also, we omit the use of $\text{CLS}]$ tokens as our task is residue-level classification. This branch facilitate the modeling of long-range interactions between residues, preserving spatial information of discontiunous epitopes.
3.2.4 Feature coupling units
To enable bidirectional interaction between the local CNN features and global Transformer representations, we introduce FCUs starting from the second block. FCUs consist of two modules: FCU-Down, which passes CNN features to the Transformer; and FCU-Up, which feeds Transformer outputs back to the CNN.
FCU-Down projects the CNN features $\mathbf{F}_{\mathrm{CNN}}\in\mathbb{R}^{C_{1}\times L}$ into the Transformer embedding space $\mathbb{R}^{L\times D}$ via a $1\times 1$ convolution and batch normalization, followed by transposition to align the dimensions. The projected features $\mathbf{F}_{\mathrm{CNN}}^{\rightarrow}$ are then added to the Transformer input $\mathbf{F}_{\mathrm{Trans}}$:
$\mathbf{F}_{\mathrm{CNN}}^{\rightarrow}$ $=\left(\mathrm{BN}\big{(}\mathrm{Conv}_{1\times 1}(\mathbf{F}_{\mathrm{CNN}})\big{)}\right)^{T},$ (6)
$\widetilde{\mathbf{F}}_{\mathrm{Trans}}$ $=\mathbf{F}_{\mathrm{Trans}}+\mathbf{F}_{\mathrm{CNN}}^{\rightarrow}.$
FCU-Up maps the Transformer features $\widetilde{\mathbf{F}}_{\mathrm{Trans}}\in\mathbb{R}^{L\times D}$ back to the CNN feature space $\mathbb{R}^{C_{2}\times L}$ by transposing first, then applying a $1\times 1$ convolution and batch normalization. The transformed features $\mathbf{F}_{\mathrm{Trans}}^{\leftarrow}$ are added to the original CNN features $\mathbf{F}_{\mathrm{CNN}}$:
$\mathbf{F}_{\mathrm{Trans}}^{\leftarrow}$ $=\mathrm{BN}\big{(}\mathrm{Conv}_{1\times 1}(\widetilde{\mathbf{F}}_{\mathrm{Trans}}^{T})\big{)},$ (7)
$\widetilde{\mathbf{F}}_{\mathrm{CNN}}$ $=\mathbf{F}_{\mathrm{CNN}}+\mathbf{F}_{\mathrm{Trans}}^{\leftarrow}.$
3.3 Performance Metric
3.3.1 AgIoU
To evaluate the overlap between the prediction and real epitopes, we use the Antigen-level Intersection over Union (AgIoU), which is defined as the ratio of the number of correctly predicted epitope residues to the total number of residues that are either labeled as epitope in the ground truth or prediction as such:
$\text{AgIoU}=\frac{\text{Res}_{\text{true}}\cap\text{Res}_{\text{pred}}}{\text{Res}_{\text{true}}\cup\text{Res}_{\text{pred}}}.$ (8)
4 Experiments
4.1 Experiments Setup
4.1.1 Dataset
We collected 1,300 Ab–Ag complex FASTA files with AlphaFold-2.3 colabfold version from AbEpiTope-1.0. Each complex was then modeled with AlphaFold2 Multimer v3, generating 30 structural predictions per complex and a total of 39,000 structures. For each complex, we selected the top-ranked model and excluded those with a pLDDT score below 0.82. Interface residues were extracted based on a distance cutoff of 4 Å between heavy atoms on the antigen and antibody chains (1). To enhance data quality and limit parameter overhead, we removed complexes containing insufficient epitope residues or antigens exceeding 1024 residues in length. Finally, the dataset contains 1,180 Ag-Ab complexes in total: 1,080 used for training BConformeR, and 100 set aside for evaluation and ablation. And the antigen chains annotated with epitope labels serve as the input.
Figure 3: Improved performance on epitope prediction. ROC-AUC and PR-AUC curves plot on the blind test set of 24 antigen records. (A) ROC-AUC curves. (B) PR-AUC curves.
Method Threshold F1-L $\uparrow$ F1-D $\uparrow$ MCC $\uparrow$ PCC $\uparrow$ AgIoU $\uparrow$
SEPPA-3.0 0.100 ± 0.020 0.606 ± 0.055 0.024 ± 0.001 0.058 ± 0.014 0.067 0.088 ± 0.005
BepiPred-3.0 0.210 ± 0.020 0.657 ± 0.032 0.034 ± 0.003 0.191 ± 0.003 0.215 0.149 ± 0.003
SEMA-1D 2.0 0.510 ± 0.030 0.683 ± 0.027 0.029 ± 0.001 0.182 ± 0.002 0.217 0.143 ± 0.002
DiscoTope-3.0 0.300 ± 0.020 0.706 ± 0.015 0.047 ± 0.003 0.227 ± 0.007 0.276 0.167 ± 0.002
BConformeR 0.300 ± 0.020 0.673 ± 0.014 0.105 ± 0.004 0.302 ± 0.003 0.347 0.212 ± 0.002
Table 2: Performance comparison of different methods on a blind test of 24 antigen records. F1-L and F1-D refer to the F1 scores calculated on linear and discontinuous epitopes, respectively; Thresholds were selected following those of baseline methods; Standard deviation calculated over a range of thresholds.
Due to the differences in training and test datasets across existing epitope prediction methods, the selection of a blind test set was important for fair evaluation. Therefore, we also extracted 24 antigen records from the test sets of SEMA-1D 2.0 and DiscoTope-3.0. At least 75% of these records were previously used as test data by each baseline method, and 100% blind for BConformeR.
Model Linear Discontinuous Params (M) MACs (G)
Precision Recall F1 Precision Recall F1
Conformer-12 0.965 0.516 0.672 0.064 0.393 0.110 302.24 309.93
Conformer-9 0.946 0.489 0.645(-4.0%) 0.055 0.344 0.095(-13.6%) 230.27 236.12
Conformer-6 0.977 0.449 0.615(-8.5%) 0.041 0.263 0.071(-35.5%) 158.30 162.31
ResNet-152 0.979 0.514 0.674(+0.3%) 0.046 0.290 0.079(-28.2%) 242.44 248.62
ResNet-101 0.961 0.468 0.629(-6.4%) 0.027 0.216 0.048(-56.4%) 179.24 183.80
ViT-12 0.969 0.477 0.639(-4.9%) 0.049 0.355 0.086(-21.8%) 230.24 234.42
ViT-9 0.958 0.455 0.617(-8.2%) 0.047 0.328 0.082(-25.5%) 173.57 176.32
Table 3: Ablation study of different models evaluated on 100 antigen records. Conformer-12 corresponds to BConformeR, and Conformer-9/6 reduces one layer of FCU-based CNN and Transformer for each stage compared to Conformer-12/9. ResNet bottlenecks adopt the same CNN blocks as used in Conformer, and ViT follows the design of Transformer branch. Models are grouped with comparable Params and MACs: Conformer-9, ResNet-152 and ViT-12; Conformer-6, ResNet-101 and ViT-9.
4.1.2 Data encodings and features
To represent antigen sequences, we used pretrained ESM-2 embeddings. Sequences were tokenized with the ESM-2 tokenizer and passed through its 33-layer transformer. We extracted the hidden states from the final layer, discarded the beginning-of-sequence (BOS) and end-of-sequence (EOS) tokens, and retained only the embeddings corresponding to the actual residues. Finally, each antigen sequence was represented as a $\mathbb{R}^{1024\times 1280}$ feature matrix, with 1024 as the padded length and 1280 as the ESM-2 embedding size.
4.2 Comparison with SOTA Methods
4.2.1 Benchmark evaluation
The performance of BConformeR was evaluated on a blind test set of 24 antigen records and compared with four SOTA methods: SEPPA-3.0, BepiPred-3.0, SEMA-1D 2.0 and DiscoTope-3.0. As shown in Table 2, BConformeR outperforms baselines in MCC (0.302), PCC (0.347), and AgIoU (0.212). Further evaluation of ROC-AUC and PR-AUC shows that BConformeR achieves scores of 0.777 and 0.313 respectively, surpassing all methods that provided residue-level probabilistic outputs (Fig. 3).
4.2.2 Enhanced prediction on conformational epitopes
To evaluate the ability of each method in predicting linear and conformational epitopes, we computed the F1 scores separately on the test set. As shown in Table 2, BConformeR achieves an F1 score of 0.105 on discontinuous epitopes, significantly higher than baseline methods that typically report these scores below 0.050. This illustrates that BConformeR is more effective at identifying non-contiguous epitope fragments.
Figure 4: Structural visualization of epitopes predicted by BConformeR vs. actual epitopes. (A-C) Comparison of predicted and actual epitopes of 7KDB. Prediction is fully covered by the ground truth. (D-F) Comparison of predicted and actual epitopes of 7DFB. Prediction partially overlaps with the ground truth. (G-I) Comparison of predicted and actual epitopes of 7C88. Real epitopes are fully covered by the prediction.
4.3 Ablation
4.3.1 Overview
To further assess the contribution of each module to linear and conformational epitope prediction, we ablated the dual-branch Conformer into standalone ResNet and ViT variants. We first reduced the depth of Conformer-12 (BConformeR) to evaluate the effectiveness of the original architecture. Then, ResNet-152 and ResNet-101, as well as ViT-12 and ViT-9, were selected based on comparable parameters (Params) and multiply-accumulate operations (MACs) to Conformer-9 and Conformer-6 respectively. The F1 score was used as the primary evaluation metric for both linear and discontinuous epitope prediction.
4.3.2 Depth reduction
To evaluate the impact of model depth, we progressively reduced the number of layers in the Conformer architecture. As shown in Table 3, Conformer-12 achieves the highest F1 scores on both linear (0.672) and discontinuous (0.110) epitopes. Reducing depth to Conformer-9 leads to a moderate drop in performance, with linear and discontinuous F1 scores decreasing by 4.0% (0.645) and 13.6% (0.095) respectively. Further reduction to Conformer-6 causes a more significant decline, especially in discontinuous epitope prediction, where the F1 score drops by 35.5% to 0.071. These demonstrate that our final model, BConformeR, achieves superior performance compared to shallower variants.
4.3.3 Conformer vs. CNN
Purely convolutional variants capture contiguous residue patterns efficiently and therefore achieve competitive scores on linear epitopes. However, their strictly local receptive field prevents them from modelling long-range residue couplings that dominate conformational epitopes. As shown in Table 3, ResNet-152 attains the highest linear F1 score (0.674) among all models, confirming its effectiveness on local patterns. However, its discontinuous F1 drops to 0.079, lower than Conformer-9 (0.095) and ViT-12 (0.086), indicating a weakness in capturing discontinuous dependencies.
By injecting a Transformer branch and exchanging information through FCUs, Conformer-9 achieves a discontinuous F1 score of 0.095, outperforming ResNet-152 (0.079), while maintaining a competitive linear F1 score of 0.645. This shows that integrating global context into convolutional layers can effectively improve discontinuous epitope prediction without sacrificing linear performance.
4.3.4 Conformer vs. ViT
Vision Transformer variants excel at aggregating sequence-wide information but lack explicit inductive bias for short-range motifs, often resulting in underfitting on linear stretches and noisy attention over densely packed regions. As shown in Table 3, ViT-9 achieves a discontinuous F1 of 0.082, surpassing ResNet-152 (0.079) and ResNet-101 (0.048). However, this comes at the cost of reduced performance on linear epitopes, where ViT-9 trails behind ResNet-101 (0.617 vs. 0.629).
Conformer mitigates this trade-off, as Conformer-9 outperforms ViT-12 on both linear (0.645 vs. 0.639) and discontinuous (0.095 vs. 0.086) F1 scores. Empirically, the hybrid design offers a more balanced solution for linear and discontinuous epitope prediction.
4.4 High-Confidence Predicted Structures
4.4.1 Predicted epitopes with high AgIoU
To visualize the prediction compared to the ground truth, we selected three antigen records from the test set with high AgIoU scores. Three overlap cases were shown in Figure 4: (1) the prediction is fully covered by the real epitopes (Fig. 4A-C), (2) the prediction partially overlaps with the real epitopes (Fig. 4D-F), (3) the real epitopes are entirely encompassed by the prediction (Fig. 4G-I).
5 Conclusion and Limitation
We present BConformeR, a dual–branch Conformer model that integrates convolutional and self–attention mechanisms for residue–level prediction of both linear and conformational B-cell epitopes from antigen sequences. In addition to the architecture itself, two auxiliary contributions are worth noting: (i) an adaptive logit–fusion mechanism that automatically balances evidence from the CNN and Transformer branches, yielding a single, well-calibrated probability for each residue without extra post-processing; (ii) a detailed performance analysis on both linear and discontinuous epitopes, which also validates the complementary effectiveness of the CNN and Transformer branches. Benchmark evaluation demonstrates that our model consistently outperforms strong sequence- and structure-based baselines, particularly on the more difficult task of predicting conformational epitopes. The anticipated beneficiaries include rational vaccine design, diagnostic reagent development and antibody engineering pipelines.
Limitations.
There are two notable limitations to this study. First, all training and test interfaces are derived from in silico AlphaFold-Multimer models; although this choice enlarges the dataset, it may introduce misoriented chains or misplaced atoms. Second, despite the size of our collection, antigen families and epitope types are imbalanced, which could bias both training and evaluation.
Acknowledgments
The computation resources in this study was sponsored by BayVax Biotech Limited. Y.-F. H. thanks support from the Major Program of the National Natural Science Foundation of China (92369201). The research was also supported by the National Key Research and Development Program of China (2021YFA0910700), the Health and Medical Research Fund, the Food and Health Bureau, The Government of the Hong Kong Special Administrative Region (COVID1903010, T-11-709/21-N) to J.-D.H. J.-D.H. also thanks the L & T Charitable Foundation, the Program for Guangdong Introducing Innovative and Entrepreneurial Teams (2019BT02Y198) and Shenzhen Key Laboratory for Cancer Metastasis and Personalized Therapy (ZDSYS20210623091811035) for their support.
Appendix
Figure 5: Related work on epitope prediction methods. FFNN denotes feedforward neural network; FC layers denotes fully connected layers; MSA denotes multiple sequence alignment results.
Figure 6: Linear regions and discontinuous fragments. (A) (Left) The blue structure represents antigen 7c88, and the light green and light orange correspond to parts of the two antibody heavy chains; Linear regions are marked with red rectangles, and discontinuous fragments with green. (Right) Residues highlighted with numbers indicate epitopes. (B) A linear region of 7c88. Gap size of up to three non-contacting residues is chosen for a linear region on the basis of the structure of $\alpha$-helices.
Appendix A Antibody-agnostic prediction methods
We summarize widely-used epitope prediction methods, categorized into sequence-based and structure-based approaches, as shown in Figure 5. Methods with the latest versions are summarized along with their data representation strategies and classifiers.
Appendix B Linear Regions and Discontinuous Fragments
Regions are defined as linear stretches of antigen sequence having at least three antibody-contacting residues (Fig. 6A). Gaps between contacting residues are allowed, and a gap size of up to three non-contacting residues is chosen on the basis of the structure of $\alpha$-helices, which enables residues on the same helical face to be included (Fig. 6B). Fragments, which do not contain enough antibody-contacting residues to qualify as a region (less than three), are considered discontinuous in our study (Fig. 6A).
Appendix C Dataset
C.1 Structure prediction
We initially obtained 1730 antibody–antigen (Ab–Ag) FASTA files from AbEpiTope-1.0. However, due to the limited computational time and the large size of certain input sequences, we only processed 1300 complexes using the AlphaFold2 Multimer v3. Each of these 1300 complexes was modeled with 5 seeds, 6 models and 3 recycles, generating 30 structures per complex.
C.2 Epitope statistics
The final dataset consists of 1180 Ab–Ag complexes, which contains a total of 1432 antigen chains and 2320 antibody chains. Epitopes were identified using a 4 Å distance cutoff between antigen and antibody residues. After classifying the epitopes into linear and discontinuous types, we found 2388 linear regions comprising 15,886 residues, and 2632 discontinuous fragments comprising 3574 residues.
Appendix D Performance Metrics
The performance metrics used in this study include Precision, Recall, F1-score, Antigen-level intersection over union (AgIoU), Matthews correlation coefficient (MCC) and Pearson correlation coefficient (PCC), whose definitions are given as follows:
$\text{Precision}=\frac{TP}{TP+FP},$ (9)
$\text{Recall}=\frac{TP}{TP+FN},$
$\text{F1}=\frac{2TP}{2TP+FP+FN},$
$\text{AgIoU}=\frac{TP}{TP+FP+FN},$
$\text{MCC}=\frac{TP\cdot TN-FP\cdot FN}{\sqrt{(TP+FP)(TP+FN)(TN+FP)(TN+FN)}},$
$\text{PCC}=\frac{\sum_{i=1}^{n}(y_{i}-\bar{y})(\hat{y}_{i}-\overline{\hat{y}})}{\sqrt{\sum_{i=1}^{n}(y_{i}-\bar{y})^{2}}\sqrt{\sum_{i=1}^{n}(\hat{y}_{i}-\overline{\hat{y}})^{2}}},$
where $TP$, TNTN, FPFP and $FN$ denote the number of true positives, true negatives, false positives and false negatives, respectively; $y_{i}$ and $\hat{y}_{i}$ represent the ground-truth and predicted values, while $\bar{y}$ and $\overline{\hat{y}}$ denote their means. The relationships between these metrics are given as:
$\text{F1}=\frac{2\cdot\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}}=\frac{2\cdot\text{AgIoU}}{1+\text{AgIoU}},$ (10)
$\text{AgIoU}=\frac{2\cdot\text{F1}}{2-\text{F1}}.$
Appendix E Evaluation and Comparison
E.1 Evaluation
Figure 7: Metrics across thresholds on the test set of 100 antigens.
BConformeR was evaluated at various thresholds on 100 antigen records. As shown in Figure 7, the range between 0.1 and 0.5—highlighted by gray dashed lines—represents the region where precision, recall, F1 score and AgIoU perform well. Notably, the F1 score peaks around 0.44 when thresholds are set between 0.1 and 0.3.
E.2 Baseline methods
Thresholds for all baselines were adjusted to reflect their best scoring behavior, ensuring a fair comparison with BConformeR in the blind test. We observed that SEPPA-3.0 performs better under a low decision threshold (around 0.1). For SEMA-1D 2.0, which defines epitope residues using an 8 Å distance cutoff, the predicted scores tend to be higher; thus, we adopted a threshold of 0.5 above its default of 0.351. Moreover, although the default threshold of BepiPred-3.0 is 0.1512, we found that a slightly higher value around 0.2 led to better balance between precision and recall on our test set.
Appendix F Ablation
F.1 ViT
Model Embed Dimension Seq Length Patch Size Depth Heads MLP Ratio
ViT-9 1536 1024 1 9 12 2.0
ViT-12 1536 1024 1 12 12 2.0
Table 4: ViT-1D variants used in ablation study.
We conducted ablation studies on two variants of the ViT-1D model differing in depth: ViT-9 and ViT-12. As shown in Table 4, both models share the same embedding dimension (1536), input sequence length (1024), patch size (1), number of attention heads (12), and MLP expansion ratio (2.0). The only difference is the number of transformer encoder layers.
F.2 Resnet
Stage Output Channels Kernel Size Resnet-101 Resnet-152
input 1280 - - -
stem 320 7 - -
layer1 640 (160 $\times$ 4) 3 3 3
layer2 1280 (320 $\times$ 4) 3 4 8
layer3 2560 (640 $\times$ 4) 3 23 36
layer4 5120 (1280 $\times$ 4) 3 3 3
classifier 2 1 - -
Table 5: ResNet-1D architectures used in ablation study.
We employed two variants of the ResNet-1D backbone in the ablation studies: ResNet-101 and ResNet-152. As shown in Table 5, both models accept input feature maps with 1280 channels and begin with an initial convolutional layer producing 320 channels. The networks consist of four residual stages of bottlenecks with the expansion factor of 4, each stage starting by downsampling the channel dimension to half. The difference lies in the number of bottleneck blocks per stage. ResNet1D-101 contains [3, 4, 23, 3] blocks across the four layers, while ResNet1D-152 is deeper with [3, 8, 36, 3] blocks. The final output is projected to the target classes through a $\times1$ convolutional layer.
Appendix G Data and Code Availability
The data and code are available at https://github.com/Jerry-alt19/BConformeR .

### v2

@@
 Ablation studies demonstrate that CNN enhances the prediction of linear epitopes, and the Transformer module improves the prediction of conformational epitopes.
-Experimental results show that our model outperforms existing baselines in terms of PCC, ROC-AUC, PR-AUC, and F1 scores on conformational epitopes.
-Keywords B-cell Epitope Prediction; Conformer; Antigen–Antibody Interface
+Experimental results show that our model outperforms existing baselines in terms of PCC, ROC-AUC, PR-AUC, and F1 scores on both linear and conformational epitopes.
+The data and code are available at https://github.com/Jerry-alt19/BConformeR.
+Keywords B-cell Epitope Prediction; Antigen–Antibody Interface; Conformer
 1 Introduction
@@
 Linear epitopes are formed by linear stretches of amino acid residues, while conformational epitopes are formed by distant residues brought together in 3D through folding.
-Figure 1: Overview of epitope classification and BConformerR prediction.
+Figure 1: Overview of epitope classification and BConformeR prediction.
 (A) Epitopes, i.e., Antibody-binding sites, include conformational fragments (n=2,632) and linear regions (n=2,388).
-Data derived from 1,432 antigen records.
-(B) Epitope prediction workflow: Antigen inputs are processed to predict linear and conformational epitopes, which enable applications in vaccine design and diagnostic development.
+Data were derived from 1,432 antigen records (Appendix C.2).
+(B) Epitope prediction workflow.
+Antigen inputs are processed to predict linear and conformational epitopes, which enable applications in vaccine design and diagnostic development.
 Identifying epitopes is critical for antibody engineering koz:18 ; ji:22 , disease diagnosis va:22 ; ta:24 , and vaccine optimization dv:24 ; ba:25 .
@@
 (A) Preparation pipeline of the data.
-1300 antigen-antibody FASTA files were processed using AlphaFold Multimer v3 to predict structures, with top-ranked structure for each was selected to extract epitopes (right); Antigen sequences were embedded by ESM-2, producing embeddings of dimension 1024×1280 (left).
+(Right) 1300 antigen-antibody FASTA files were processed using AlphaFold Multimer v3 to predict structures, with top-ranked structure for each was selected to extract epitopes.
+(Left) Antigen sequences were embedded by ESM-2, producing embeddings of dimension 1024×1280.
 (B) Conformer architecture.
@@
 To address this multi-scale complexity, we designed the Conformer architecture, which integrates convolution layers to learn local residue-level features and self-attention mechanisms to capture long-range dependencies within antigen sequences Conformer20 ; Conformer21 .
-Here, we propose BConformeR, a Conformer-based model for linear and conformational BCR epitope prediction.
+Here, we propose BConformeR, a Conformer-based model for continuous and discontinuous BCR epitope prediction.
 Antigen sequences from 1,180 Ag-Ab complexes were embedded using ESM-2, with 1,080 used for training and 100 for evaluation.
-In a blind test of 24 antigen records, BConformeR outperforms existing baselines—SEPPA-3.0, BepiPred-3.0, SEMA-1D 2.0 and DiscoTope-3.0—across multiple metrics, including PCC, ROC-AUC, PR-AUC, and F1 score on conformational epitopes.
+In a blind test of 24 antigen records, BConformeR outperforms existing baselines—SEPPA-3.0, BepiPred-3.0, SEMA-1D 2.0 and DiscoTope-3.0—across multiple metrics, including PCC, ROC-AUC, PR-AUC, and F1 score on both linear and conformational epitopes.
 To further dissect the model’s design, we ablated ResNet bottlenecks and ViT modules from Conformer, showing that convolutional layers improve linear epitope prediction, and Transformer modules are critical for capturing discontinuous epitope patterns.
@@
 2.1 Antibody-agnostic prediction models
-Numerous models, trained on antigen–epitope pairs, have been proposed and refined to enhance epitope prediction.
+Numerous models, trained on antigen–epitope pairs, have been proposed and refined to enhance epitope prediction (Appendix A).
 SEPPA-3.0 was the epitope prediction tool specifically designed for glycoproteins, based on the logistic regression classifier SEPPA3.0 .
@@
 In this work, we utilize the real Ag-Ab complex dataset provided by AbEpitope-1.0, employing its epitope identification pipeline to ensure the data quality (Fig. 2A).
-3 Method
+3 Methods
 3.1 Epitope Extraction
@@
 Regions are defined as linear stretches of antigen sequence having at least three antibody-contacting residues.
-Gaps between contacting residues are allowed, and a gap size of up to three non-contacting residues is chosen on the basis of the structure of $\alpha$-helices, which enables residues on the same helical face to be included epitope_cl .
+Gaps between contacting residues are allowed, and a gap size of up to three non-contacting residues is chosen on the basis of the structure of $\alpha$-helices, which enables residues on the same helical face to be included epitope_cl (Appendix B).
 Fragments, which do not contain enough antibody-contacting residues to qualify as a region, are considered discontinuous in our study (Fig.
@@
 $\alpha\times\mathrm{Conv}_{\mathrm{cls}}+(1-\alpha)\times\mathrm{Trans}_{\mathrm{cls}}$
-Table 1: Architecture of BConformeR, a 1D adaptation of Conformer Conformer21 .
+Table 1: Architecture of BConformeR (BConformer-12), a 1D adaptation of Conformer Conformer21 .
 The MHSA-12 indicates multi-head self-attention with 12 heads operating on 1536-dimensional embeddings (128 dimensions per head).
@@
 $\mathcal{F}(\mathbf{x})=\begin{cases}\mathbf{x}.&\text{if shape matches}\\ \text{BN}_{\text{res}}\left(\text{Conv}_{1\times 1}(\mathbf{x})\right).&\text{otherwise}\end{cases}$
-The channel dimension is first downsampled to half, then expanded by a factor of 4 within each bottleneck.
+The channel dimension is first downsampled, then expanded by a factor of 4 within each bottleneck.
 The first convolution block contains a single bottleneck, while subsequent blocks include two bottlenecks: one applied before the Transformer branch, and the other afterward.
@@
 $\mathbf{z}_{2}$ $=\mathbf{z}_{1}+\mathrm{DropPath}\left(\mathrm{MLP}\left(\mathrm{LN}_{2}(\mathbf{z}_{1})\right)\right),$
-where $\mathrm{MHSA}(\cdot)$ denotes the multi-head self-attention mechanism with $h$ (e.g., 12) heads, and $\mathrm{MLP}(\cdot)$ consists of a linear up-projection to $2D$ dimensions, a GELU activation, followed by a linear down-projection to $D$.
+where $\mathrm{MHSA}(\cdot)$ denotes the multi-head self-attention mechanism with $h$ (e.g., 12) heads Attention , and $\mathrm{MLP}(\cdot)$ consists of a linear up-projection to $2D$ dimensions, a GELU activation, followed by a linear down-projection to $D$.
 The $\mathrm{DropPath}$ operator implements stochastic depth regularization.
@@
 3.3.1 AgIoU
-To evaluate the overlap between the prediction and real epitopes, we use the Antigen-level Intersection over Union (AgIoU), which is defined as the ratio of the number of correctly predicted epitope residues to the total number of residues that are either labeled as epitope in the ground truth or prediction as such:
+To evaluate the overlap between the prediction and real epitopes, we use the Antigen-level Intersection over Union (AgIoU), which is defined as the ratio of the number of correctly predicted epitope residues to the total number of residues that are either labeled as epitope in the ground truth or prediction as such (Appendix D):
 $\text{AgIoU}=\frac{\text{Res}_{\text{true}}\cap\text{Res}_{\text{pred}}}{\text{Res}_{\text{true}}\cup\text{Res}_{\text{pred}}}.$ (8)
@@
 4.1.1 Dataset
-We collected 1,300 Ab–Ag complex FASTA files with AlphaFold-2.3 colabfold version from AbEpiTope-1.0.
-Each complex was then modeled with AlphaFold2 Multimer v3, generating 30 structural predictions per complex and a total of 39,000 structures.
+We collected 1,300 Ag–Ab complex FASTA files with AlphaFold-2.3 colabfold version from AbEpiTope-1.0.
+Each complex was then modeled with AlphaFold2 Multimer v3, generating 30 structural predictions per complex and a total of 39,000 structures (Appendix C.1).
 For each complex, we selected the top-ranked model and excluded those with a pLDDT score below 0.82.
@@
 Method Threshold F1-L $\uparrow$ F1-D $\uparrow$ MCC $\uparrow$ PCC $\uparrow$ AgIoU $\uparrow$
-SEPPA-3.0 0.100 ± 0.020 0.606 ± 0.055 0.024 ± 0.001 0.058 ± 0.014 0.067 0.088 ± 0.005
-BepiPred-3.0 0.210 ± 0.020 0.657 ± 0.032 0.034 ± 0.003 0.191 ± 0.003 0.215 0.149 ± 0.003
-SEMA-1D 2.0 0.510 ± 0.030 0.683 ± 0.027 0.029 ± 0.001 0.182 ± 0.002 0.217 0.143 ± 0.002
-DiscoTope-3.0 0.300 ± 0.020 0.706 ± 0.015 0.047 ± 0.003 0.227 ± 0.007 0.276 0.167 ± 0.002
-BConformeR 0.300 ± 0.020 0.673 ± 0.014 0.105 ± 0.004 0.302 ± 0.003 0.347 0.212 ± 0.002
+SEPPA-3.0 0.100 ± 0.020 0.158 ± 0.009 0.237 ± 0.025 0.058 ± 0.014 0.067 0.088 ± 0.005
+BepiPred-3.0 0.220 ± 0.010 0.276 ± 0.007 0.209 ± 0.011 0.194 ± 0.003 0.215 0.155 ± 0.003
+SEMA-1D 2.0 0.500 ± 0.030 0.270 ± 0.005 0.213 ± 0.006 0.188 ± 0.002 0.217 0.147 ± 0.002
+DiscoTope-3.0 0.300 ± 0.020 0.303 ± 0.007 0.231 ± 0.004 0.227 ± 0.007 0.276 0.167 ± 0.002
+BConformeR 0.330 ± 0.020 0.349 ± 0.004 0.262 ± 0.003 0.296 ± 0.003 0.347 0.203 ± 0.002
 Table 2: Performance comparison of different methods on a blind test of 24 antigen records.
-F1-L and F1-D refer to the F1 scores calculated on linear and discontinuous epitopes, respectively; Thresholds were selected following those of baseline methods; Standard deviation calculated over a range of thresholds.
+F1-L and F1-D refer to the F1 scores calculated on linear and discontinuous epitopes, respectively; Thresholds were selected following those of baseline methods (Appendix E.2); Standard deviation calculated over a range of thresholds.
 Due to the differences in training and test datasets across existing epitope prediction methods, the selection of a blind test set was important for fair evaluation.
-Therefore, we also extracted 24 antigen records from the test sets of SEMA-1D 2.0 and DiscoTope-3.0.
-At least 75% of these records were previously used as test data by each baseline method, and 100% blind for BConformeR.
+Therefore, we also extracted 24 antigen records from the test sets of SEMA-1D 2.0 and DiscoTope-3.0, all of which are blind to BConformeR.
 Model Linear Discontinuous Params (M) MACs (G)
 Precision Recall F1 Precision Recall F1
-Conformer-12 0.965 0.516 0.672 0.064 0.393 0.110 302.24 309.93
-Conformer-9 0.946 0.489 0.645(-4.0%) 0.055 0.344 0.095(-13.6%) 230.27 236.12
-Conformer-6 0.977 0.449 0.615(-8.5%) 0.041 0.263 0.071(-35.5%) 158.30 162.31
-ResNet-152 0.979 0.514 0.674(+0.3%) 0.046 0.290 0.079(-28.2%) 242.44 248.62
-ResNet-101 0.961 0.468 0.629(-6.4%) 0.027 0.216 0.048(-56.4%) 179.24 183.80
-ViT-12 0.969 0.477 0.639(-4.9%) 0.049 0.355 0.086(-21.8%) 230.24 234.42
-ViT-9 0.958 0.455 0.617(-8.2%) 0.047 0.328 0.082(-25.5%) 173.57 176.32
-Table 3: Ablation study of different models evaluated on 100 antigen records.
-Conformer-12 corresponds to BConformeR, and Conformer-9/6 reduces one layer of FCU-based CNN and Transformer for each stage compared to Conformer-12/9.
-ResNet bottlenecks adopt the same CNN blocks as used in Conformer, and ViT follows the design of Transformer branch.
-Models are grouped with comparable Params and MACs: Conformer-9, ResNet-152 and ViT-12; Conformer-6, ResNet-101 and ViT-9.
+BConformer-12 0.440 0.477 0.458 0.304 0.360 0.330 302.24 309.93
+BConformer-9 0.428 0.463 0.445(-2.8%) 0.297 0.322 0.309(-6.4%) 230.27 236.12
+BConformer-6 0.423 0.416 0.419(-8.5%) 0.249 0.296 0.270(-18.2%) 158.30 162.31
+ResNet1D-152 0.442 0.484 0.462(+0.9%) 0.255 0.291 0.272(-17.6%) 242.44 248.62
+ResNet1D-101 0.417 0.430 0.423(-7.6%) 0.242 0.238 0.240(-27.3%) 179.24 183.80
+ViT1D-12 0.429 0.435 0.432(-5.7%) 0.269 0.334 0.298(-9.7%) 230.24 234.42
+ViT1D-9 0.403 0.428 0.415(-9.4%) 0.276 0.311 0.292(-11.5%) 173.57 176.32
+Table 3: Ablation study of different models evaluated on 100 antigen records (Appendix E.1).
+BConformer-12 corresponds to BConformeR, and BConformer-9/6 reduces one layer of FCU-based CNN and Transformer for each stage compared to BConformer-12/9.
+ResNet bottlenecks adopt the same CNN blocks as used in Conformer, and ViT follows the design of Transformer branch (Appendix F).
+Models are grouped with comparable Params and MACs: BConformer-9, ResNet1D-152 and ViT1D-12; BConformer-6, ResNet1D-101 and ViT1D-9.
 4.1.2 Data encodings and features
@@
 3).
-4.2.2 Enhanced prediction on conformational epitopes
-To evaluate the ability of each method in predicting linear and conformational epitopes, we computed the F1 scores separately on the test set.
-As shown in Table 2, BConformeR achieves an F1 score of 0.105 on discontinuous epitopes, significantly higher than baseline methods that typically report these scores below 0.050.
-This illustrates that BConformeR is more effective at identifying non-contiguous epitope fragments.
+4.2.2 Enhanced prediction on continuous and discontinuous epitopes
+To evaluate the ability of each method in predicting continuous and discontinuous epitopes, we computed the F1 scores separately on the test set.
+As shown in Table 2, BConformeR achieves F1 scores of 0.349 and 0.262 on linear and discontinuous epitopes, respectively, higher than all baseline methods.
+This illustrates that BConformeR is more effective at identifying both regions and fragments of epitopes.
 Figure 4: Structural visualization of epitopes predicted by BConformeR vs.
@@
 To further assess the contribution of each module to linear and conformational epitope prediction, we ablated the dual-branch Conformer into standalone ResNet and ViT variants.
-We first reduced the depth of Conformer-12 (BConformeR) to evaluate the effectiveness of the original architecture.
-Then, ResNet-152 and ResNet-101, as well as ViT-12 and ViT-9, were selected based on comparable parameters (Params) and multiply-accumulate operations (MACs) to Conformer-9 and Conformer-6 respectively.
+We first reduced the depth of BConformer-12 (BConformeR) to evaluate the effectiveness of the original architecture.
+Then, ResNet1D-152 and ResNet1D-101, as well as ViT1D-12 and ViT1D-9, were selected based on comparable parameters (Params) and multiply-accumulate operations (MACs) to BConformer-9 and BConformer-6 respectively.
 The F1 score was used as the primary evaluation metric for both linear and discontinuous epitope prediction.
@@
 To evaluate the impact of model depth, we progressively reduced the number of layers in the Conformer architecture.
-As shown in Table 3, Conformer-12 achieves the highest F1 scores on both linear (0.672) and discontinuous (0.110) epitopes.
-Reducing depth to Conformer-9 leads to a moderate drop in performance, with linear and discontinuous F1 scores decreasing by 4.0% (0.645) and 13.6% (0.095) respectively.
-Further reduction to Conformer-6 causes a more significant decline, especially in discontinuous epitope prediction, where the F1 score drops by 35.5% to 0.071.
+As shown in Table 3, reducing depth to BConformer-9 leads to a moderate drop in performance, with linear and discontinuous F1 scores decreasing by 2.8% (0.445) and 6.4% (0.309) respectively.
+Further reduction to BConformer-6 causes a more significant decline, especially in discontinuous epitope prediction, where the F1 score drops by 18.2% to 0.270.
 These demonstrate that our final model, BConformeR, achieves superior performance compared to shallower variants.
@@
 However, their strictly local receptive field prevents them from modelling long-range residue couplings that dominate conformational epitopes.
-As shown in Table 3, ResNet-152 attains the highest linear F1 score (0.674) among all models, confirming its effectiveness on local patterns.
-However, its discontinuous F1 drops to 0.079, lower than Conformer-9 (0.095) and ViT-12 (0.086), indicating a weakness in capturing discontinuous dependencies.
-By injecting a Transformer branch and exchanging information through FCUs, Conformer-9 achieves a discontinuous F1 score of 0.095, outperforming ResNet-152 (0.079), while maintaining a competitive linear F1 score of 0.645.
-This shows that integrating global context into convolutional layers can effectively improve discontinuous epitope prediction without sacrificing linear performance.
+As shown in Table 3, ResNet1D-152 attains the highest linear F1 score (0.462) among all models, confirming its effectiveness on local patterns.
+However, its discontinuous F1 drops to 0.272, lower than BConformer-9 (0.309) and ViT1D-12 (0.298), pointing to difficulty in capturing discontinuous dependencies.
+By injecting a Transformer branch and exchanging information through FCUs, BConformer-9 achieves a noticeably higher discontinuous F1 than ResNet1D-152, while keeping a linear F1 of 0.445 with only a minor drop.
+This shows that substituting part of the convolutional layers with a Transformer modules improves discontinuous epitope prediction with little sacrifice in linear performance.
 4.3.4 Conformer vs. ViT
 Vision Transformer variants excel at aggregating sequence-wide information but lack explicit inductive bias for short-range motifs, often resulting in underfitting on linear stretches and noisy attention over densely packed regions.
-As shown in Table 3, ViT-9 achieves a discontinuous F1 of 0.082, surpassing ResNet-152 (0.079) and ResNet-101 (0.048).
-However, this comes at the cost of reduced performance on linear epitopes, where ViT-9 trails behind ResNet-101 (0.617 vs.
-0.629).
-Conformer mitigates this trade-off, as Conformer-9 outperforms ViT-12 on both linear (0.645 vs.
-0.639) and discontinuous (0.095 vs.
-0.086) F1 scores.
+As shown in Table 3, ViT1D-9 achieves a discontinuous F1 of 0.292, surpassing ResNet1D-101 (0.240), ResNet1D-152 (0.272) and BConformer-6 (0.270).
+However, this comes at the cost of reduced performance on linear epitopes, where ViT1D-9 (0.415) trails behind ResNet1D-101 (0.423) and BConformer-6 (0.419).
+Conformer mitigates this trade-off, as BConformer-9 outperforms ViT1D-12 on both linear and discontinuous F1 scores.
 Empirically, the hybrid design offers a more balanced solution for linear and discontinuous epitope prediction.
@@
 Limitations.
-There are two notable limitations to this study.
+There are three notable limitations to this study.
 First, all training and test interfaces are derived from in silico AlphaFold-Multimer models; although this choice enlarges the dataset, it may introduce misoriented chains or misplaced atoms.
-Second, despite the size of our collection, antigen families and epitope types are imbalanced, which could bias both training and evaluation.
+Second, antigen families and epitope types remain imbalanced, which could bias both training and evaluation.
+This is reflected in the significantly higher F1 scores observed on the 100-antigen test set compared with the 24-antigen blind test set, even though all records are unseen by BConformeR.
+Third, considering the computational cost, input sequences are limited to 1024 residues.
+Antigens longer than this threshold (e.g., 2nyy, 3ri5) cannot be directly processed by the model.
 Acknowledgments
@@
 Figure 6: Linear regions and discontinuous fragments.
-(A) (Left) The blue structure represents antigen 7c88, and the light green and light orange correspond to parts of the two antibody heavy chains; Linear regions are marked with red rectangles, and discontinuous fragments with green.
+(A) (Left) The blue structure represents antigen 7c88, and the light green and light orange correspond to part of the antibody light chain and heavy chain, respectively; Linear regions are marked with red rectangles, and discontinuous fragments with green.
 (Right) Residues highlighted with numbers indicate epitopes.
@@
 6A).
+An algorithm classifying linear epitopes for an antigen sequence is shown below:
+Algorithm 1 Linear Epitope Classification
+1:label_tensorlabel\_tensor : sequence of residue labels
+2:   ($1=$ contacting, $0=$ non-contacting, 100=-100= padding)
+3:linear_flagslinear\_flags: boolean mask indicating linear epitope residues
+4:Convert label_tensorlabel\_tensor to list labelslabels of length $L$
+5:Initialize $linear\_flags\leftarrow[\text{False}]^{L}$, $visited\leftarrow[\text{False}]^{L}$
+[…]
+7:while $i<L$ do
+8:  if $labels[i]\neq 1$ or $visited[i]=$ True then
+9:   $i\leftarrow i+1$ $\triangleright$ skip non-epitope or already visited residue
+10:   continue
+11:  end if
+12:  Initialize $segment\_indices\leftarrow[\,]$, $one\_count\leftarrow 0$, $zero\_count\leftarrow 0$
+13:  $j\leftarrow i$
+14:  while $j<L$ do
+15:   if labels[j]=100labels[j]=-100 then
+16:     break $\triangleright$ stop at padding
+17:   else if $labels[j]=1$ then
+18:     Append $j$ to segment_indicessegment\_indices
+19:     $one\_count\leftarrow one\_count+1$, $zero\_count\leftarrow 0$
+20:   else if $labels[j]=0$ then
+21:     $zero\_count\leftarrow zero\_count+1$
+22:     if zero_count>3zero\_count>3 then
+23:      break $\triangleright$ gap too large
+24:     end if
+25:     Append $j$ to segment_indicessegment\_indices
+26:   end if
+27:   $j\leftarrow j+1$
+28:  end while
+29:  $ones\_in\_segment\leftarrow\{k\in segment\_indices\mid labels[k]=1\}$
+30:  if $|ones\_in\_segment|\geq 3$ then
+31:   for $k\in ones\_in\_segment$ do
+32:     $linear\_flags[k]\leftarrow\text{True}$, $visited[k]\leftarrow\text{True}$
+33:   end for
+34:   $i\leftarrow segment\_indices[-1]+1$
+35:  else
+36:   $visited[i]\leftarrow\text{True}$, $i\leftarrow i+1$
+37:  end if
+38:end while
+39:return linear_flagslinear\_flags
 Appendix C Dataset
 C.1 Structure prediction
-We initially obtained 1730 antibody–antigen (Ab–Ag) FASTA files from AbEpiTope-1.0.
+We initially obtained 1730 Ag-Ab FASTA files from AbEpiTope-1.0.
 However, due to the limited computational time and the large size of certain input sequences, we only processed 1300 complexes using the AlphaFold2 Multimer v3.
@@
 C.2 Epitope statistics
-The final dataset consists of 1180 Ab–Ag complexes, which contains a total of 1432 antigen chains and 2320 antibody chains.
+The final dataset consists of 1180 Ag–Ab complexes, which contains a total of 1432 antigen chains and 2360 antibody chains.
 Epitopes were identified using a 4 Å distance cutoff between antigen and antibody residues.
@@
 Appendix D Performance Metrics
-The performance metrics used in this study include Precision, Recall, F1-score, Antigen-level intersection over union (AgIoU), Matthews correlation coefficient (MCC) and Pearson correlation coefficient (PCC), whose definitions are given as follows:
-$\text{Precision}=\frac{TP}{TP+FP},$ (9)
+The performance metrics used in this study include Precision, Recall, True Positive Rate (TPR), False Positive Rate (FPR), F1-score, Antigen-level intersection over union (AgIoU), Matthews correlation coefficient (MCC) and Pearson correlation coefficient (PCC), whose definitions are given as follows:
+$\text{Precision}=\frac{TP}{TP+FP},$
 $\text{Recall}=\frac{TP}{TP+FN},$
+$\text{TPR}=\frac{TP}{TP+FN},$
+$\text{FPR}=\frac{FP}{FP+TN},$
 $\text{F1}=\frac{2TP}{2TP+FP+FN},$
@@
 where $TP$, TNTN, FPFP and $FN$ denote the number of true positives, true negatives, false positives and false negatives, respectively; $y_{i}$ and $\hat{y}_{i}$ represent the ground-truth and predicted values, while $\bar{y}$ and $\overline{\hat{y}}$ denote their means.
-The relationships between these metrics are given as:
-$\text{F1}=\frac{2\cdot\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}}=\frac{2\cdot\text{AgIoU}}{1+\text{AgIoU}},$ (10)
+Note that Recall and TPR are mathematically equivalent, but used in different contexts.
+The relationships between some of these metrics are given as:
+$\text{F1}=\frac{2\cdot\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}}=\frac{2\cdot\text{AgIoU}}{1+\text{AgIoU}},$
 $\text{AgIoU}=\frac{2\cdot\text{F1}}{2-\text{F1}}.$
@@
 For SEMA-1D 2.0, which defines epitope residues using an 8 Å distance cutoff, the predicted scores tend to be higher; thus, we adopted a threshold of 0.5 above its default of 0.351.
-Moreover, although the default threshold of BepiPred-3.0 is 0.1512, we found that a slightly higher value around 0.2 led to better balance between precision and recall on our test set.
+Moreover, although the default threshold of BepiPred-3.0 is 0.1512, we found that a slightly higher value around 0.22 led to better balance between precision and recall on our test set.
 Appendix F Ablation
-F.1 ViT
+F.1 ViT1D
 Model Embed Dimension Seq Length Patch Size Depth Heads MLP Ratio
-ViT-9 1536 1024 1 9 12 2.0
-ViT-12 1536 1024 1 12 12 2.0
-Table 4: ViT-1D variants used in ablation study.
-We conducted ablation studies on two variants of the ViT-1D model differing in depth: ViT-9 and ViT-12.
+ViT1D-9 1536 1024 1 9 12 2.0
+ViT1D-12 1536 1024 1 12 12 2.0
+Table 4: ViT variants used in ablation study.
+We conducted ablation studies on two variants of the ViT1D model differing in depth: ViT1D-9 and ViT1D-12.
 As shown in Table 4, both models share the same embedding dimension (1536), input sequence length (1024), patch size (1), number of attention heads (12), and MLP expansion ratio (2.0).
 The only difference is the number of transformer encoder layers.
-F.2 Resnet
-Stage Output Channels Kernel Size Resnet-101 Resnet-152
-input 1280 - - -
-stem 320 7 - -
+F.2 ResNet1D
+Stage Output Channels Kernel Size ResNet1D-101 ResNet1D-152
+input 1280 - $\checkmark$ $\checkmark$
+stem 320 7 $\checkmark$ $\checkmark$
 layer1 640 (160 $\times$ 4) 3 3 3
@@
 layer4 5120 (1280 $\times$ 4) 3 3 3
-classifier 2 1 - -
-Table 5: ResNet-1D architectures used in ablation study.
-We employed two variants of the ResNet-1D backbone in the ablation studies: ResNet-101 and ResNet-152.
+classifier 2 1 $\checkmark$ $\checkmark$
+Table 5: ResNet architectures used in ablation study.
+We employed two variants of the ResNet1D backbone in the ablation studies: ResNet1D-101 and ResNet1D-152.
 As shown in Table 5, both models accept input feature maps with 1280 channels and begin with an initial convolutional layer producing 320 channels.
@@
 The final output is projected to the target classes through a $\times1$ convolutional layer.
-Appendix G Data and Code Availability
-The data and code are available at https://github.com/Jerry-alt19/BConformeR .

### v3

@@
 Abstract
-Accurate prediction of antibody-binding sites (epitopes) on antigens is crucial for vaccine design, immunodiagnostics, therapeutic antibody development, antibody engineering, research into autoimmune and allergic diseases, and for advancing our understanding of immune responses.
+Accurate prediction of antibody-binding sites (epitopes) on antigens is crucial for vaccine design, immunodiagnostics, therapeutic antibody development, antibody engineering, research into autoimmune and allergic diseases, and advancing our understanding of immune responses.
 Despite in silico methods that have been proposed to predict both linear (continuous) and conformational (discontinuous) epitopes, they consistently underperform in predicting conformational epitopes.
-In this work, we propose a conformer-based model trained on antigen sequences derived from 1,080 antigen-antibody complexes, leveraging convolutional neural networks (CNNs) to extract local features and Transformers to capture long-range dependencies within antigen sequences.
+In this work, we propose Conformer-based models trained separately on AlphaFold-predicted structures and experimentally determined structures, leveraging convolutional neural networks (CNNs) to extract local features and Transformers to capture long-range dependencies within antigen sequences.
 Ablation studies demonstrate that CNN enhances the prediction of linear epitopes, and the Transformer module improves the prediction of conformational epitopes.
-Experimental results show that our model outperforms existing baselines in terms of PCC, ROC-AUC, PR-AUC, and F1 scores on both linear and conformational epitopes.
-The data and code are available at https://github.com/Jerry-alt19/BConformeR.
+Experimental results show that our model outperforms existing baselines in terms of MCC, ROC-AUC, PR-AUC, and F1 scores on both linear and conformational epitopes.
+Code is available at https://anonymous.4open.science/r/bconformer-epitope-519D.
 Keywords B-cell Epitope Prediction; Antigen–Antibody Interface; Conformer
@@
 (A) Epitopes, i.e., Antibody-binding sites, include conformational fragments (n=2,632) and linear regions (n=2,388).
-Data were derived from 1,432 antigen records (Appendix C.2).
+Data were derived from 1,432 antigen records.
 (B) Epitope prediction workflow.
-Antigen inputs are processed to predict linear and conformational epitopes, which enable applications in vaccine design and diagnostic development.
+Antigen inputs are processed to predict linear and conformational epitopes, which enables applications in vaccine design and diagnostic development.
 Identifying epitopes is critical for antibody engineering koz:18; ji:22, disease diagnosis va:22; ta:24, and vaccine optimization dv:24; ba:25.
@@
 Figure 2: BConformeR overview.
-(A) Preparation pipeline of the data.
-(Right) 1300 antigen-antibody FASTA files were processed using AlphaFold Multimer v3 to predict structures, with top-ranked structure for each was selected to extract epitopes.
-(Left) Antigen sequences were embedded by ESM-2, producing embeddings of dimension 1024×1280.
-(B) Conformer architecture.
-Stem module applies convolution and max-pooling to the original ESM-2 embeddings; The Conv block and Trans block are based on ResNet bottleneck and Vision Transformer modules respectively; Feature coupling units (FCU-Up and FCU-Down) connect the Conv and Trans blocks, with 11 FCU layers in total; The classifier linearly fuses features from both Conv and Trans branches.
+(A) Preparation pipeline.
+(Right) Preparation of Alphafold-predicted structures and epitope extraction.
+(Left) Antigen embeddings.
+(B) Architecture overview.
+Stem module applies convolution and max-pooling to the ESM Cambrian (ESMC) embeddings; The Conv block and Trans block are based on ResNet bottleneck and Vision Transformer modules respectively; FCU-Up and FCU-Down connect the Conv and Trans blocks; The classifier linearly fuses features from both branches.
 However, epitope prediction remains a challenging task, with existing methods underperform on conformational ones.
@@
 Here, we propose BConformeR, a Conformer-based model for continuous and discontinuous BCR epitope prediction.
-Antigen sequences from 1,180 Ag-Ab complexes were embedded using ESM-2, with 1,080 used for training and 100 for evaluation.
-In a blind test of 24 antigen records, BConformeR outperforms existing baselines—SEPPA-3.0, BepiPred-3.0, SEMA-1D 2.0 and DiscoTope-3.0—across multiple metrics, including PCC, ROC-AUC, PR-AUC, and F1 score on both linear and conformational epitopes.
+BConformeR outperforms baselines—SEPPA-3.0, BepiPred-3.0, SEMA-1D 2.0 and DiscoTope-3.0—across multiple metrics, including MCC, ROC-AUC, PR-AUC, and F1 score on both linear and conformational epitopes.
 To further dissect the model’s design, we ablated ResNet bottlenecks and ViT modules from Conformer, showing that convolutional layers improve linear epitope prediction, and Transformer modules are critical for capturing discontinuous epitope patterns.
-2 Related Work
-2.1 Antibody-agnostic prediction models
-Numerous models, trained on antigen–epitope pairs, have been proposed and refined to enhance epitope prediction (Appendix A).
-SEPPA-3.0 was the epitope prediction tool specifically designed for glycoproteins, based on the logistic regression classifier SEPPA3.0 .
-BepiPred-3.0 used embeddings from pretrained protein language models ESM-2 as input to a feedforward neural network (FFNN) to predict both linear and conformational B-cell epitopes BepiPred3.0 .
-Epitope3D modeled antigen surfaces as residue-level graphs and extracted distance-based graph signatures to train classifiers for conformational B-cell epitope prediction, with AdaBoost selected as the backbone Epitope3D ; AdaBoost .
-ScanNet was an end-to-end geometric deep learning model that learned interpretable residue-level features directly from 3D structural neighborhoods ScanNet .
-SEMA-1D 2.0 was constructed by adding a fully-connected linear layer on an ensemble of five ESM-2 models, while SEMA-3D 2.0 was developed using an ensemble of five pre-trained Structure-aware Protein language models (SaProt) SEMA2.0 ; SaProt .
-DiscoTope-3.0 utilized inverse folding representations from ESM-IF1 and was trained on both predicted and solved structures using a positive-unlabelled ensemble strategy, enabling structure-based B-cell epitope prediction at scale DiscoTope3.0 ; ESM-IF1 .
-However, existing models rarely incorporate both local features and long-range dependencies within antigen sequences, showing limited performance on conformational epitopes.
-Therefore, we adopted a Conformer based on mutual sampling, employing CNNs to extract local features and Transformers to capture long-range dependencies within antigen sequences.
-2.2 Antibody-aware evaluation models
-Despite advances in epitope prediction, evaluating Ag-Ab binding confidence remains challenging, which is crucial for vaccine selection.
-AbEpitope-1.0, a recent method for predicting antibody targets by use of AlphaFold and inverse folding, significantly improving the prediction accuracy of Ag-Ab interfaces.
-They constructed negative Ag-Ab samples (false binding pairs) by swapping antibody chains across real complexes in the dataset.
-Each sample was labeled as either true or false to train the model to distinguish authentic Ag-Ab interactions from decoys.
-The final model, AbEpiTarget-1.0, achieved superior performance in antibody-sensitive scoring for target prediction AbEpitope .
-In this work, we utilize the real Ag-Ab complex dataset provided by AbEpitope-1.0, employing its epitope identification pipeline to ensure the data quality (Fig. 2A).
 2 Methods
-3.1 Epitope Extraction
-3.1.1 Epitope identification
-During immune recognition, epitopes are typically composed of multiple spatially adjacent residues.
-To capture this interaction, the notions of residue-neighbor and residue-patch were introduced.
-A residue-neighbor is defined when the minimum distance between any atoms of two residues is less than 4 Å, and a residue-patch refers to a group of residues whose atoms lie within 10 Å of a central residue epitope_id .
-In our study, we focused on the residue-neighbor relationship to capture the tight spatial interactions, which is given as:
-$\min_{a\in r_{i},\,b\in r_{j}}\|a-b\|<4\,\text{\AA }$, (1)
-where $r_{i}$ and $r_{j}$ are residues from the antigen and antibody chains respectively, and $a$, $b$ represent the heavy atoms within these residues.
-3.1.2 Epitope classification
+2.1 Epitope classification
+To extract epitope residues, we define the residue–neighbor, where an antigen residue and an antibody residue are considered binding sites if the minimum distance between any pair of their heavy atoms is less than 4 Å epitope_id.
 Generally, epitopes are separated into multiple Regions (R) and Fragments (F) for each antigen sequence.
 Regions are defined as linear stretches of antigen sequence having at least three antibody-contacting residues.
-Gaps between contacting residues are allowed, and a gap size of up to three non-contacting residues is chosen on the basis of the structure of $\alpha$-helices, which enables residues on the same helical face to be included epitope_cl (Appendix B).
+Gaps between contacting residues are allowed, and a gap size of up to three non-contacting residues is chosen on the basis of the structure of $\alpha$-helices, which enables residues on the same helical face to be included epitope_cl.
 Fragments, which do not contain enough antibody-contacting residues to qualify as a region, are considered discontinuous in our study (Fig.
@@
 where $N$, $L$, and $D$ denote non-epitope, linear epitope and discontinuous epitope, respectively. Performance metrics (e.g., F1) are then calculated for each type of epitopes.
-stage output CNN Branch FCU Transformer Branch
-stem 1024 7, 320
-3 max pooling
-s1 1024 $\left[\begin{array}[]{l}\text{1, 160}\\[-1.00006pt] \text{3, 160}\\[-1.00006pt] \text{1, 640}\end{array}\right]$ - 5, 1536 $\times$1
-$\left[\begin{array}[]{c}\text{\text{MHSA-12}, 1536}\\[-1.00006pt] \text{1, 3072}\\[-1.00006pt] \text{1, 1536}\end{array}\right]$
-$\left[\begin{array}[]{l}\text{1, 160}\\[-1.00006pt] \text{3, 160}\\[-1.00006pt] \text{1, 640}\end{array}\right]$ $\times$3
-$1,1536]\longrightarrow$
-$\left[\begin{array}[]{c}\text{\text{MHSA-12}, 1536}\\[-1.00006pt] \text{1, 3072}\\[-1.00006pt] \text{1, 1536}\end{array}\right]$
-$\left[\begin{array}[]{l}\text{1, 160}\\[-1.00006pt] \text{3, 160}\\[-1.00006pt] \text{1, 640}\end{array}\right]$
-$\longleftarrow[1,160]$
-s2 1024 $\left[\begin{array}[]{l}\text{1, 320}\\[-1.00006pt] \text{3, 320}\\[-1.00006pt] \text{1, 1280}\end{array}\right]$ $\times$4
-$1,1536]\longrightarrow$
-$\left[\begin{array}[]{c}\text{\text{MHSA-12}, 1536}\\[-1.00006pt] \text{1, 3072}\\[-1.00006pt] \text{1, 1536}\end{array}\right]$
-$\left[\begin{array}[]{l}\text{1, 320}\\[-1.00006pt] \text{3, 320}\\[-1.00006pt] \text{1, 1280}\end{array}\right]$
-$\longleftarrow[1,320]$
-s3 1024 $\left[\begin{array}[]{l}\text{1, 640}\\[-1.00006pt] \text{3, 640}\\[-1.00006pt] \text{1, 2560}\end{array}\right]$ $\times$4
-$1,1536]\longrightarrow$
-$\left[\begin{array}[]{c}\text{\text{MHSA-12}, 1536}\\[-1.00006pt] \text{1, 3072}\\[-1.00006pt] \text{1, 1536}\end{array}\right]$
-$\left[\begin{array}[]{l}\text{1, 640}\\[-1.00006pt] \text{3, 640}\\[-1.00006pt] \text{1, 2560}\end{array}\right]$
-$\longleftarrow[1,640]$
-[…]
-$\alpha\times\mathrm{Conv}_{\mathrm{cls}}+(1-\alpha)\times\mathrm{Trans}_{\mathrm{cls}}$
-Table 1: Architecture of BConformeR (BConformer-12), a 1D adaptation of Conformer Conformer21 .
-The MHSA-12 indicates multi-head self-attention with 12 heads operating on 1536-dimensional embeddings (128 dimensions per head).
-$\mathrm{Conv}_{\mathrm{cls}}$ denotes the 1D convolutional classifier operating on CNN features, and $\mathrm{Trans}_{\mathrm{cls}}$ denotes the linear classifier operating on Transformer outputs.
-The final prediction is obtained by a weighted fusion of the two.
+  stage 1 stage 2 stage 3   CNN $\left[\begin{array}[]{l}1,144\\[-1.99997pt] 3,144\\[-1.99997pt] 1,576\end{array}\right]$ $\left[\begin{array}[]{l}1,144\\[-1.99997pt] 3,144\\[-1.99997pt] 1,576\end{array}\right]$ $\left[\begin{array}[]{l}1,144\\[-1.99997pt] 3,144\\[-1.99997pt] 1,576\end{array}\right]$ $\left[\begin{array}[]{l}1,288\\[-1.99997pt] 3,288\\[-1.99997pt] 1,1152\end{array}\right]$ $\left[\begin{array}[]{l}1,288\\[-1.99997pt] 3,288\\[-1.99997pt] 1,1152\end{array}\right]$ $\left[\begin{array}[]{l}1,576\\[-1.99997pt] 3,576\\[-1.99997pt] 1,2304\end{array}\right]$ $\left[\begin{array}[]{l}1,576\\[-1.99997pt] 3,576\\[-1.99997pt] 1,2304\end{array}\right]$   FCU - $\begin{array}[]{c}[1,1536]\\[-1.99997pt]$ $\big\downarrow\end{array}$ $\begin{array}[]{c}\big\uparrow\\[-1.99997pt] [1,144]\end{array}$ $\begin{array}[]{c}[1,1536]\\[-1.99997pt]$ $\big\downarrow\end{array}$ $\begin{array}[]{c}\big\uparrow\\[-1.99997pt] [1,288]\end{array}$ $\begin{array}[]{c}[1,1536]\\[-1.99997pt]$ $\big\downarrow\end{array}$ $\begin{array}[]{c}\big\uparrow\\[-1.99997pt] [1,576]\end{array}$   Trans $\left[\begin{array}[]{c}\text{MHSA-24, 1536}\\[-1.99997pt] 1,3072\\[-1.99997pt] 1,1536\end{array}\right]$ $\left[\begin{array}[]{c}\text{MHSA-24, 1536}\\[-1.99997pt] 1,3072\\[-1.99997pt] 1,1536\end{array}\right]$ $\left[\begin{array}[]{c}\text{MHSA-24, 1536}\\[-1.99997pt] 1,3072\\[-1.99997pt] 1,1536\end{array}\right]$ $\left[\begin{array}[]{c}\text{MHSA-24, 1536}\\[-1.99997pt] 1,3072\\[-1.99997pt] 1,1536\end{array}\right]$   $\times 1$ $\times 3$ $\times 4$ $\times 4$
+Table 1: Interaction layers of BConformeR. MHSA denotes multi-head self-attention.
 2.2 Conformer
 2.2.1 Overview
-To take advantage of both local and global representations for epitope prediction, we construct BConformeR, a dual-branch architecture adapted from Conformer Conformer21 , as shown in Table 1.
+To take advantage of both local and global representations for epitope prediction, we construct BConformeR, a dual-branch architecture adapted from Conformer Conformer21.
 BConformeR consists of a shared convolutional stem, a CNN branch, a Transformer branch, Feature Coupling Units (FCUs) for bidirectional feature exchange, and two per-position classifiers (Fig.
 2B).
-Since our task is per-residue epitope prediction, all antigen sequences are padded to a fixed length of $L$ (e.g., 1024) and this length is preserved throughout the model.
-The stem consists of a 1D convolution (kernel size 7) and a 1D max pooling (kernel size 3), designed to capture contextual features for each residue.
-And the embedding dimension $C$ (e.g., 1280) is reduced to $C/4$ channels after the stem.
-In the first block, both branches process identical initial features and therefore no feature coupling is applied.
+Sequences are padded to match the longest length in each batch.
+The stem consists of a 1D convolution and a 1D max pooling, with the embedding dimension $C$ (e.g., 1152 generated by ESMC-600M) reduced to $C/4$ channels after the stem.
+Then as shown in Table 1, both branches process initial features and no feature coupling is applied in the first block.
 From the second block onward, FCUs are inserted to progressively fuse local and global representations.
-The input is processed by alternating CNN and Transformer blocks, where at each stage, CNN-derived local features are passed to the Transformer branch, and Transformer-derived global representations are fed back to the CNN branch.
+At each stage, CNN-derived features are passed to the Transformer branch, and Transformer-derived global representations are fed back to the CNN branch.
 The CNN branch gradually expands to $2C$ channels, and the Transformer branch maintains $D$-dimensional (e.g., 1536) token embeddings at the end.
@@
 The CNN branch is composed of three stages after the stem, each consisting of repeated convolutional blocks based on the ResNet bottleneck design ResNet.
-Given an input feature sequence $\mathbf{x}\in\mathbb{R}^{C^{\prime}\times L}$, a bottleneck performs three sequential convolutions with a residual connection applied to $\mathbf{x}$:
+As shown in Table 1, given an input $\mathbf{x}\in\mathbb{R}^{C^{\prime}\times L}$, the bottleneck performs three sequential convolutions with a residual connection applied to $\mathbf{x}$:
 $\mathbf{z}_{1}$ $=\sigma\left(\text{BN}_{1}\left(\text{Conv}_{1\times 1}(\mathbf{x})\right)\right),$ (3)
@@
 The channel dimension is first downsampled, then expanded by a factor of 4 within each bottleneck.
-The first convolution block contains a single bottleneck, while subsequent blocks include two bottlenecks: one applied before the Transformer branch, and the other afterward.
-Since convolution kernels slide over representation sequences with overlap, the CNN blocks are well-suited to capture residue-level contextual information, providing local features for the Transformer branch.
+The first convolution block contains a single bottleneck, while subsequent blocks include bottlenecks before and after the Transformer branch.
 2.2.3 Transformer branch
-The Transformer branch processes residue embeddings using 12 repeated Transformer blocks.
-Each block follows the standard architecture introduced in ViT ViT , but adapted to 1D protein sequences.
-As shown in Table 1, each block contains a multi-head self-attention module and a two-layer feed-forward MLP block.
-Both submodules are preceded by Layer Normalization (LN) with residual connections.
-Specifically, given the input sequence $\mathbf{x}\in\mathbb{R}^{L\times D}$, a single Transformer block updates the representation as follows:
+The Transformer branch processes embeddings using 12 repeated Transformer blocks, with each block following a structure similar to ViT ViT.
+As shown in Table 1, given the input $\mathbf{x}\in\mathbb{R}^{L\times D}$, a single Transformer block updates the representation as follows:
 $\mathbf{z}_{1}$ $=\mathbf{x}+\mathrm{DropPath}\left(\mathrm{MHSA}\left(\mathrm{LN}_{1}(\mathbf{x})\right)\right),$ (4)
 $\mathbf{z}_{2}$ $=\mathbf{z}_{1}+\mathrm{DropPath}\left(\mathrm{MLP}\left(\mathrm{LN}_{2}(\mathbf{z}_{1})\right)\right),$
-where $\mathrm{MHSA}(\cdot)$ denotes the multi-head self-attention mechanism with $h$ (e.g., 12) heads Attention , and $\mathrm{MLP}(\cdot)$ consists of a linear up-projection to $2D$ dimensions, a GELU activation, followed by a linear down-projection to $D$.
+where $\mathrm{MHSA}(\cdot)$ denotes the multi-head self-attention mechanism with $h$ (e.g., 24) heads Attention, and $\mathrm{MLP}(\cdot)$ consists of a linear up-projection to $2D$ dimensions, a GELU activation, followed by a linear down-projection to $D$.
 The $\mathrm{DropPath}$ operator implements stochastic depth regularization.
@@
 2.2.4 Feature coupling units
-To enable bidirectional interaction between the local CNN features and global Transformer representations, we introduce FCUs starting from the second block.
-FCUs consist of two modules: FCU-Down, which passes CNN features to the Transformer; and FCU-Up, which feeds Transformer outputs back to the CNN.
-FCU-Down projects the CNN features $\mathbf{F}_{\mathrm{CNN}}\in\mathbb{R}^{C_{1}\times L}$ into the Transformer embedding space $\mathbb{R}^{L\times D}$ via a $1\times 1$ convolution and batch normalization, followed by transposition to align the dimensions.
-The projected features $\mathbf{F}_{\mathrm{CNN}}^{\rightarrow}$ are then added to the Transformer input $\mathbf{F}_{\mathrm{Trans}}$:
+To enable bidirectional interaction between local CNN features and global Transformer representations, we introduce feature coupling units (FCUs) from the second block.
+FCU-Down projects CNN features into the Transformer space and adds them to the Transformer input:
 $\mathbf{F}_{\mathrm{CNN}}^{\rightarrow}$ $=\left(\mathrm{BN}\big(\mathrm{Conv}_{1\times 1}(\mathbf{F}_{\mathrm{CNN}})\big)\right)^{T},$ (5)
 $\widetilde{\mathbf{F}}_{\mathrm{Trans}}$ $=\mathbf{F}_{\mathrm{Trans}}+\mathbf{F}_{\mathrm{CNN}}^{\rightarrow}.$
-FCU-Up maps the Transformer features $\widetilde{\mathbf{F}}_{\mathrm{Trans}}\in\mathbb{R}^{L\times D}$ back to the CNN feature space $\mathbb{R}^{C_{2}\times L}$ by transposing first, then applying a $1\times 1$ convolution and batch normalization.
-The transformed features $\mathbf{F}_{\mathrm{Trans}}^{\leftarrow}$ are added to the original CNN features $\mathbf{F}_{\mathrm{CNN}}$:
+FCU-Up maps Transformer features back to the CNN space and merges them:
 $\mathbf{F}_{\mathrm{Trans}}^{\leftarrow}$ $=\mathrm{BN}\big(\mathrm{Conv}_{1\times 1}(\widetilde{\mathbf{F}}_{\mathrm{Trans}}^{T})\big),$ (6)
 $\widetilde{\mathbf{F}}_{\mathrm{CNN}}$ $=\mathbf{F}_{\mathrm{CNN}}+\mathbf{F}_{\mathrm{Trans}}^{\leftarrow}.$
-3.3 Performance Metric
-3.3.1 AgIoU
-To evaluate the overlap between the prediction and real epitopes, we use the Antigen-level Intersection over Union (AgIoU), which is defined as the ratio of the number of correctly predicted epitope residues to the total number of residues that are either labeled as epitope in the ground truth or prediction as such (Appendix D):
-$\text{AgIoU}=\frac{\text{Res}_{\text{true}}\cap\text{Res}_{\text{pred}}}{\text{Res}_{\text{true}}\cup\text{Res}_{\text{pred}}}.$ (8)
+2.3 Calibrated Score
+Scores of residues in each antigen are normalized using the following formula:
+$z_{i}=\frac{s_{i}-\mu_{\text{antigen}}}{\sigma_{\text{antigen}}},$ (7)
+where $s_{i}$ denotes the raw score of residue $i$, and $\mu_{\text{antigen}}$ and $\sigma_{\text{antigen}}$ represent the mean and standard deviation of scores for that antigen, respectively.
+Both $\mu_{\text{antigen}}$ and $\sigma_{\text{antigen}}$ are estimated using two separate generalized additive models (GAMs) GAM, with one predicting the mean residue score from antigen length and the other predicting the standard deviation from the mean residue score DiscoTope3.0.
 3 Experiments
@@
 3.1.1 Dataset
-We collected 1,300 Ag–Ab complex FASTA files with AlphaFold-2.3 colabfold version from AbEpiTope-1.0.
-Each complex was then modeled with AlphaFold2 Multimer v3, generating 30 structural predictions per complex and a total of 39,000 structures (Appendix C.1).
-For each complex, we selected the top-ranked model and excluded those with a pLDDT score below 0.82.
-Interface residues were extracted based on a distance cutoff of 4 Å between heavy atoms on the antigen and antibody chains (1).
-To enhance data quality and limit parameter overhead, we removed complexes containing insufficient epitope residues or antigens exceeding 1024 residues in length.
-Finally, the dataset contains 1,180 Ag-Ab complexes in total: 1,080 used for training BConformeR, and 100 set aside for evaluation and ablation.
-And the antigen chains annotated with epitope labels serve as the input.
-Figure 3: Improved performance on epitope prediction. ROC-AUC and PR-AUC curves plot on the blind test set of 24 antigen records. (A) ROC-AUC curves. (B) PR-AUC curves.
-Method Threshold F1-L $\uparrow$ F1-D $\uparrow$ MCC $\uparrow$ PCC $\uparrow$ AgIoU $\uparrow$
-SEPPA-3.0 0.100 ± 0.020 0.158 ± 0.009 0.237 ± 0.025 0.058 ± 0.014 0.067 0.088 ± 0.005
-BepiPred-3.0 0.220 ± 0.010 0.276 ± 0.007 0.209 ± 0.011 0.194 ± 0.003 0.215 0.155 ± 0.003
-SEMA-1D 2.0 0.500 ± 0.030 0.270 ± 0.005 0.213 ± 0.006 0.188 ± 0.002 0.217 0.147 ± 0.002
-DiscoTope-3.0 0.300 ± 0.020 0.303 ± 0.007 0.231 ± 0.004 0.227 ± 0.007 0.276 0.167 ± 0.002
-BConformeR 0.330 ± 0.020 0.349 ± 0.004 0.262 ± 0.003 0.296 ± 0.003 0.347 0.203 ± 0.002
-Table 2: Performance comparison of different methods on a blind test of 24 antigen records.
-F1-L and F1-D refer to the F1 scores calculated on linear and discontinuous epitopes, respectively; Thresholds were selected following those of baseline methods (Appendix E.2); Standard deviation calculated over a range of thresholds.
-Due to the differences in training and test datasets across existing epitope prediction methods, the selection of a blind test set was important for fair evaluation.
-Therefore, we also extracted 24 antigen records from the test sets of SEMA-1D 2.0 and DiscoTope-3.0, all of which are blind to BConformeR.
-Model Linear Discontinuous Params (M) MACs (G)
-Precision Recall F1 Precision Recall F1
-BConformer-12 0.440 0.477 0.458 0.304 0.360 0.330 302.24 309.93
-BConformer-9 0.428 0.463 0.445(-2.8%) 0.297 0.322 0.309(-6.4%) 230.27 236.12
-BConformer-6 0.423 0.416 0.419(-8.5%) 0.249 0.296 0.270(-18.2%) 158.30 162.31
-ResNet1D-152 0.442 0.484 0.462(+0.9%) 0.255 0.291 0.272(-17.6%) 242.44 248.62
-ResNet1D-101 0.417 0.430 0.423(-7.6%) 0.242 0.238 0.240(-27.3%) 179.24 183.80
-ViT1D-12 0.429 0.435 0.432(-5.7%) 0.269 0.334 0.298(-9.7%) 230.24 234.42
-ViT1D-9 0.403 0.428 0.415(-9.4%) 0.276 0.311 0.292(-11.5%) 173.57 176.32
-Table 3: Ablation study of different models evaluated on 100 antigen records (Appendix E.1).
-BConformer-12 corresponds to BConformeR, and BConformer-9/6 reduces one layer of FCU-based CNN and Transformer for each stage compared to BConformer-12/9.
-ResNet bottlenecks adopt the same CNN blocks as used in Conformer, and ViT follows the design of Transformer branch (Appendix F).
-Models are grouped with comparable Params and MACs: BConformer-9, ResNet1D-152 and ViT1D-12; BConformer-6, ResNet1D-101 and ViT1D-9.
-4.1.2 Data encodings and features
-To represent antigen sequences, we used pretrained ESM-2 embeddings.
-Sequences were tokenized with the ESM-2 tokenizer and passed through its 33-layer transformer.
-We extracted the hidden states from the final layer, discarded the beginning-of-sequence (BOS) and end-of-sequence (EOS) tokens, and retained only the embeddings corresponding to the actual residues.
-Finally, each antigen sequence was represented as a $\mathbb{R}^{1024\times 1280}$ feature matrix, with 1024 as the padded length and 1280 as the ESM-2 embedding size.
+We trained two models separately using experimentally solved and AlphaFold-predicted structures.
+The dataset of experimentally solved Ag–Ab complexes was obtained from AACDB AACDB_data.
+Redundant complexes were removed by retaining only one complex for each PDB ID, resulting in a curated dataset of 3,674 Ag-Ab complexes for training.
+For the AlphaFold-based dataset, we collected 1,300 Ag–Ab complex FASTA files from AbEpiTope-1.0 AbEpitope and modeled each complex using AlphaFold-2.3 ColabFold AFMultimer.
+Thirty structural predictions were generated for each complex, producing a total of 39,000 structures.
+For each complex, we retained the top-ranked model and discarded those with a pLDDT score below 0.82.
+After extracting epitopes for each complex, we further excluded those with insufficient epitope residues.
+The final dataset contained 1,180 Ag–Ab complexes, of which 1,080 were used for training BConformeR and 100 for validation and ablation studies.
+Figure 3: Improved performance of BConformeR.
+(A–B) GAMs for estimating $\mu$ and std.
+(C–D) Comparison between raw and calibrated scores on the SARS-CoV-2 test set.
+(E–F) ROC-AUC and PR-AUC; BConformeR trained on AlphaFold-predicted structures and evaluated on the blind test set of 24 antigens.
+(G) PR-AUC; BConformeR trained on experimentally solved structures and evaluated on the SARS-CoV-2 test set.
+As the datasets used among methods are not consistent, we assembled a SARS-CoV-2 test set from CoV-AbDab CoV-AbDab for models trained on experimentally solved structures, and selected 24 antigen records from the SEMA-1D 2.0 test sets to construct a blind test set for AlphaFold-based models.
+3.1.2 Calibrated score
+After training, the $\mu_{\text{antigen}}$ and $\sigma_{\text{antigen}}$ were estimated from the predicted score distributions (Fig.
+3A,B).
+The scores were then calibrated based on different antigen lengths, which effectively reduced false positives while improving the true positive rate (Fig.
+3C,D).
+Method Alphafold structures Solved structures
+F1-L $\uparrow$ F1-D $\uparrow$ MCC $\uparrow$ F1-L $\uparrow$ F1-D $\uparrow$ MCC $\uparrow$
+SEPPA-3.0 0.158 ± 0.009 0.237 ± 0.025 0.058 ± 0.014 - - -
+BepiPred-3.0 0.276 ± 0.007 0.209 ± 0.011 0.194 ± 0.003 0.299 ± 0.009 0.087 ± 0.001 0.153 ± 0.005
+SEMA-1D 2.0 0.270 ± 0.005 0.213 ± 0.006 0.188 ± 0.002 0.327 ± 0.008 0.106 ± 0.007 0.193 ± 0.004
+DiscoTope-3.0 0.303 ± 0.007 0.231 ± 0.004 0.227 ± 0.007 0.396 ± 0.023 0.095 ± 0.007 0.307 ± 0.006
+BConformeR 0.349 ± 0.004 0.262 ± 0.003 0.296 ± 0.003 0.405 ± 0.004 0.135 ± 0.014 0.350 ± 0.014
+Table 2: Performance comparison of different methods.
+F1-L and F1-D refer to the F1 scores calculated on linear and discontinuous epitopes, respectively.
+Standard deviation calculated over a range of thresholds.
 3.2 Comparison with SOTA Methods
 3.2.1 Benchmark evaluation
-The performance of BConformeR was evaluated on a blind test set of 24 antigen records and compared with four SOTA methods: SEPPA-3.0, BepiPred-3.0, SEMA-1D 2.0 and DiscoTope-3.0.
-As shown in Table 2, BConformeR outperforms baselines in MCC (0.296), PCC (0.347), and AgIoU (0.203).
-Further evaluation of ROC-AUC and PR-AUC shows that BConformeR achieves scores of 0.777 and 0.313 respectively, surpassing all methods that provided residue-level probabilistic outputs (Fig.
-3).
-4.2.2 Enhanced prediction on continuous and discontinuous epitopes
-To evaluate the ability of each method in predicting continuous and discontinuous epitopes, we computed the F1 scores separately on the test set.
-As shown in Table 2, BConformeR achieves F1 scores of 0.349 and 0.262 on linear and discontinuous epitopes, respectively, higher than all baseline methods.
-This illustrates that BConformeR is more effective at identifying both regions and fragments of epitopes.
+BConformeR outperforms state-of-the-art baselines, achieving superior ROC-AUC and PR-AUC on Alphafold-predicted structures (Fig.
+3E,F) and higher PR-AUC on SARS-CoV-2 test set (Fig.
+3G).
+Notably, it consistently improves the prediction of both linear and discontinuous epitopes (Tab.
+2), demonstrating robust identification of epitope regions and fragments.
+Model Linear Discontinuous Params (M) MACs (G)
+Prec $\uparrow$ Rec $\uparrow$ F1 $\uparrow$ Prec $\uparrow$ Rec $\uparrow$ F1 $\uparrow$
+BConformer-12 0.440 0.477 0.458 0.304 0.360 0.330 289.31 296.64
+BConformer-9 0.428 0.463 0.445(-2.8%) 0.297 0.322 0.309(-6.4%) 219.94 225.50
+BConformer-6 0.423 0.416 0.419(-8.5%) 0.249 0.296 0.270(-18.2%) 150.56 154.37
+ResNet1D-152 0.442 0.484 0.462(+0.9%) 0.255 0.291 0.272(-17.6%) 196.41 201.45
+ResNet1D-101 0.417 0.430 0.423(-7.6%) 0.242 0.238 0.240(-27.3%) 145.21 148.92
+ViT1D-12 0.429 0.435 0.432(-5.7%) 0.269 0.334 0.298(-9.7%) 230.05 234.22
+ViT1D-9 0.403 0.428 0.415(-9.4%) 0.276 0.311 0.292(-11.5%) 173.37 176.12
+Table 3: Ablation studies evaluated on the validation set.
+BConformer-12 corresponds to BConformeR, and BConformer-9/6 reduces one layer of FCU-based CNN and Transformer for each stage compared to BConformer-12/9.
+ResNet bottlenecks adopt the same CNN blocks as Conformer, and ViT follows the design of Transformer branch.
+3.3 Ablation
+3.3.1 Overview
+To further assess the contribution of each module to linear and conformational epitope prediction, we ablated the dual-branch Conformer into standalone ResNet and ViT variants.
+We first reduced the depth of BConformer-12 (BConformeR) to evaluate the effectiveness of the original architecture.
+Then, ResNet1D-152 and ResNet1D-101, as well as ViT1D-12 and ViT1D-9, were selected based on comparable parameters (Params) and multiply-accumulate operations (MACs) to BConformer-9 and BConformer-6 respectively.
+The F1 score was used as the primary evaluation metric for both linear and discontinuous epitope prediction.
 Figure 4: Structural visualization of epitopes predicted by BConformeR vs.
@@
 Real epitopes are fully covered by the prediction.
-4.3 Ablation
-4.3.1 Overview
-To further assess the contribution of each module to linear and conformational epitope prediction, we ablated the dual-branch Conformer into standalone ResNet and ViT variants.
-We first reduced the depth of BConformer-12 (BConformeR) to evaluate the effectiveness of the original architecture.
-Then, ResNet1D-152 and ResNet1D-101, as well as ViT1D-12 and ViT1D-9, were selected based on comparable parameters (Params) and multiply-accumulate operations (MACs) to BConformer-9 and BConformer-6 respectively.
-The F1 score was used as the primary evaluation metric for both linear and discontinuous epitope prediction.
-4.3.2 Depth reduction
-To evaluate the impact of model depth, we progressively reduced the number of layers in the Conformer architecture.
-As shown in Table 3, reducing depth to BConformer-9 leads to a moderate drop in performance, with linear and discontinuous F1 scores decreasing by 2.8% (0.445) and 6.4% (0.309) respectively.
-Further reduction to BConformer-6 causes a more significant decline, especially in discontinuous epitope prediction, where the F1 score drops by 18.2% to 0.270.
-These demonstrate that our final model, BConformeR, achieves superior performance compared to shallower variants.
 3.3.2 Conformer vs. CNN
 Purely convolutional variants capture contiguous residue patterns efficiently and therefore achieve competitive scores on linear epitopes.
-However, their strictly local receptive field prevents them from modelling long-range residue couplings that dominate conformational epitopes.
+However, their strictly local receptive field prevents them from modelling long‐range residue couplings that dominate conformational epitopes.
 As shown in Table 3, ResNet1D-152 attains the highest linear F1 score (0.462) among all models, confirming its effectiveness on local patterns.
@@
 Empirically, the hybrid design offers a more balanced solution for linear and discontinuous epitope prediction.
-4.4 High-Confidence Predicted Structures
+3.4 Case Studies
 3.4.1 Predicted epitopes with high AgIoU
@@
 4G-I).
-5 Conclusion and Limitation
+4 Related Work
+Numerous antibody-agnostic epitope prediction methods have been proposed.
+SEPPA-3.0 was the epitope prediction tool specifically designed for glycoproteins, based on the logistic regression classifier SEPPA3.0.
+BepiPred-3.0 used embeddings from pretrained protein language models ESM-2 as input to a feedforward neural network (FFNN) to predict both linear and conformational B-cell epitopes BepiPred3.0.
+Epitope3D modeled antigen surfaces as residue-level graphs and extracted distance-based graph signatures to train classifiers for conformational B-cell epitope prediction, with AdaBoost selected as the backbone Epitope3D; AdaBoost.
+ScanNet was an end-to-end geometric deep learning model that learned interpretable residue-level features directly from 3D structural neighborhoods ScanNet.
+SEMA-1D 2.0 was constructed by adding a fully-connected linear layer on an ensemble of five ESM-2 models, while SEMA-3D 2.0 was developed upon an ensemble of five pre-trained SaProt models SEMA2.0; SaProt.
+DiscoTope-3.0 utilized inverse folding representations from ESM-IF1 and was trained on both predicted and solved structures using a positive-unlabelled ensemble strategy, enabling structure-based B-cell epitope prediction at scale DiscoTope3.0; ESM-IF1.
+5 Conclusion
 We present BConformeR, a dual–branch Conformer model that integrates convolutional and self–attention mechanisms for residue–level prediction of both linear and conformational B-cell epitopes from antigen sequences.
-In addition to the architecture itself, two auxiliary contributions are worth noting: (i) an adaptive logit–fusion mechanism that automatically balances evidence from the CNN and Transformer branches, yielding a single, well-calibrated probability for each residue without extra post-processing; (ii) a detailed performance analysis on both linear and discontinuous epitopes, which also validates the complementary effectiveness of the CNN and Transformer branches.
+Two contributions are worth noting: (i) Adaptive feature coupling and logit fusion that balance information from the CNN and Transformer branches, yielding a well-calibrated score for each residue; (ii) a detailed performance analysis on both linear and discontinuous epitopes, which also validates the complementary effectiveness of both branches.
 Benchmark evaluation demonstrates that our model consistently outperforms strong sequence- and structure-based baselines, particularly on the more difficult task of predicting conformational epitopes.
 The anticipated beneficiaries include rational vaccine design, diagnostic reagent development and antibody engineering pipelines.
-Limitations.
-There are three notable limitations to this study.
-First, all training and test interfaces are derived from in silico AlphaFold-Multimer models; although this choice enlarges the dataset, it may introduce misoriented chains or misplaced atoms.
-Second, antigen families and epitope types remain imbalanced, which could bias both training and evaluation.
-This is reflected in the significantly higher F1 scores observed on the 100-antigen test set compared with the 24-antigen blind test set, even though all records are unseen by BConformeR.
-Third, considering the computational cost, input sequences are limited to 1024 residues.
-Antigens longer than this threshold (e.g., 2nyy, 3ri5) cannot be directly processed by the model.
 Acknowledgments
@@
 Appendix
-Figure 5: Related work on epitope prediction methods. FFNN denotes feedforward neural network; FC layers denotes fully connected layers; MSA denotes multiple sequence alignment results.
 Figure 5: Linear regions and discontinuous fragments.
-(A) (Left) The blue structure represents antigen 7c88, and the light green and light orange correspond to part of the antibody light chain and heavy chain, respectively; Linear regions are marked with red rectangles, and discontinuous fragments with green.
+(Left) The blue structure represents antigen 7c88, and the light green and light orange correspond to part of the antibody light chain and heavy chain, respectively; Linear regions are marked with red rectangles, and discontinuous fragments with green.
 (Right) Residues highlighted with numbers indicate epitopes.
-(B) A linear region of 7c88.
-Gap size of up to three non-contacting residues is chosen for a linear region on the basis of the structure of $\alpha$-helices.
-Appendix A Antibody-agnostic prediction methods
-We summarize widely-used epitope prediction methods, categorized into sequence-based and structure-based approaches, as shown in Figure 5.
-Methods with the latest versions are summarized along with their data representation strategies and classifiers.
-Appendix B Linear Regions and Discontinuous Fragments
-Regions are defined as linear stretches of antigen sequence having at least three antibody-contacting residues (Fig.
-6A).
-Gaps between contacting residues are allowed, and a gap size of up to three non-contacting residues is chosen on the basis of the structure of $\alpha$-helices, which enables residues on the same helical face to be included (Fig.
-6B).
+Appendix A Linear Regions and Discontinuous Fragments
+Regions are defined as linear stretches of antigen sequence having at least three antibody-contacting residues.
+Gaps between contacting residues are allowed, and a gap size of up to three non-contacting residues is chosen on the basis of the structure of $\alpha$-helices, which enables residues on the same helical face to be included.
 Fragments, which do not contain enough antibody-contacting residues to qualify as a region (less than three), are considered discontinuous in our study (Fig.
-6A).
-An algorithm classifying linear epitopes for an antigen sequence is shown below:
-Algorithm 1 Linear Epitope Classification
-1:label_tensorlabel\_tensor : sequence of residue labels
-2:   ($1=$ contacting, $0=$ non-contacting, 100=-100= padding)
-3:linear_flagslinear\_flags: boolean mask indicating linear epitope residues
-4:Convert label_tensorlabel\_tensor to list labelslabels of length $L$
-5:Initialize $linear\_flags\leftarrow[\text{False}]^{L}$, $visited\leftarrow[\text{False}]^{L}$
-[…]
-7:while $i<L$ do
-8:  if $labels[i]\neq 1$ or $visited[i]=$ True then
-9:   $i\leftarrow i+1$ $\triangleright$ skip non-epitope or already visited residue
-10:   continue
-11:  end if
-12:  Initialize $segment\_indices\leftarrow[\,]$, $one\_count\leftarrow 0$, $zero\_count\leftarrow 0$
-13:  $j\leftarrow i$
-14:  while $j<L$ do
-15:   if labels[j]=100labels[j]=-100 then
-16:     break $\triangleright$ stop at padding
-17:   else if $labels[j]=1$ then
-18:     Append $j$ to segment_indicessegment\_indices
-19:     $one\_count\leftarrow one\_count+1$, $zero\_count\leftarrow 0$
-20:   else if $labels[j]=0$ then
-21:     $zero\_count\leftarrow zero\_count+1$
-22:     if zero_count>3zero\_count>3 then
-23:      break $\triangleright$ gap too large
-24:     end if
-25:     Append $j$ to segment_indicessegment\_indices
-26:   end if
-27:   $j\leftarrow j+1$
-28:  end while
-29:  $ones\_in\_segment\leftarrow\{k\in segment\_indices\mid labels[k]=1\}$
-30:  if $|ones\_in\_segment|\geq 3$ then
-31:   for $k\in ones\_in\_segment$ do
-32:     $linear\_flags[k]\leftarrow\text{True}$, $visited[k]\leftarrow\text{True}$
-33:   end for
-34:   $i\leftarrow segment\_indices[-1]+1$
-35:  else
-36:   $visited[i]\leftarrow\text{True}$, $i\leftarrow i+1$
-37:  end if
-38:end while
-39:return linear_flagslinear\_flags
-Appendix C Dataset
-C.1 Structure prediction
-We initially obtained 1730 Ag-Ab FASTA files from AbEpiTope-1.0.
-However, due to the limited computational time and the large size of certain input sequences, we only processed 1300 complexes using the AlphaFold2 Multimer v3.
-Each of these 1300 complexes was modeled with 5 seeds, 6 models and 3 recycles, generating 30 structures per complex.
-C.2 Epitope statistics
-The final dataset consists of 1180 Ag–Ab complexes, which contains a total of 1432 antigen chains and 2360 antibody chains.
-Epitopes were identified using a 4 Å distance cutoff between antigen and antibody residues.
-After classifying the epitopes into linear and discontinuous types, we found 2388 linear regions comprising 15,886 residues, and 2632 discontinuous fragments comprising 3574 residues.
-Appendix D Performance Metrics
-The performance metrics used in this study include Precision, Recall, True Positive Rate (TPR), False Positive Rate (FPR), F1-score, Antigen-level intersection over union (AgIoU), Matthews correlation coefficient (MCC) and Pearson correlation coefficient (PCC), whose definitions are given as follows:
-$\text{Precision}=\frac{TP}{TP+FP},$
-$\text{Recall}=\frac{TP}{TP+FN},$
-$\text{TPR}=\frac{TP}{TP+FN},$
-$\text{FPR}=\frac{FP}{FP+TN},$
-$\text{F1}=\frac{2TP}{2TP+FP+FN},$
-$\text{AgIoU}=\frac{TP}{TP+FP+FN},$
-$\text{MCC}=\frac{TP\cdot TN-FP\cdot FN}{\sqrt{(TP+FP)(TP+FN)(TN+FP)(TN+FN)}},$
-$\text{PCC}=\frac{\sum_{i=1}^{n}(y_{i}-\bar{y})(\hat{y}_{i}-\overline{\hat{y}})}{\sqrt{\sum_{i=1}^{n}(y_{i}-\bar{y})^{2}}\sqrt{\sum_{i=1}^{n}(\hat{y}_{i}-\overline{\hat{y}})^{2}}},$
-where $TP$, TNTN, FPFP and $FN$ denote the number of true positives, true negatives, false positives and false negatives, respectively; $y_{i}$ and $\hat{y}_{i}$ represent the ground-truth and predicted values, while $\bar{y}$ and $\overline{\hat{y}}$ denote their means.
-Note that Recall and TPR are mathematically equivalent, but used in different contexts.
-The relationships between some of these metrics are given as:
-$\text{F1}=\frac{2\cdot\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}}=\frac{2\cdot\text{AgIoU}}{1+\text{AgIoU}},$
+5).
+Appendix B Performance Metrics
+The performance metrics used in this study include Precision (Prec), Recall (Rec), True Positive Rate (TPR), False Positive Rate (FPR), F1-score, Antigen-level intersection over union (AgIoU), Matthews correlation coefficient (MCC), whose definitions are given as follows:
+$\text{Prec}=\frac{\text{TP}}{\text{TP}+\text{FP}},$
+$\text{Rec}=\frac{\text{TP}}{\text{TP}+\text{FN}},$
+$\text{TPR}=\frac{\text{TP}}{\text{TP}+\text{FN}},$
+$\text{FPR}=\frac{\text{FP}}{\text{FP}+\text{TN}},$
+$\text{F1}=\frac{2\text{TP}}{2\text{TP}+\text{FP}+\text{FN}},$
+$\text{AgIoU}=\frac{\text{TP}}{\text{TP}+\text{FP}+\text{FN}},$
+$\text{MCC}=\frac{\text{TP}\cdot\text{TN}-\text{FP}\cdot\text{FN}}{\sqrt{(\text{TP}+\text{FP})(\text{TP}+\text{FN})(\text{TN}+\text{FP})(\text{TN}+\text{FN})}},$
+where TP, TN, FP and FN denote the number of true positives, true negatives, false positives and false negatives, respectively. The relationships between some of these metrics are given as:
+$\text{F1}=\frac{2\cdot\text{Prec}\cdot\text{Rec}}{\text{Prec}+\text{Rec}}=\frac{2\cdot\text{AgIoU}}{1+\text{AgIoU}},$
 $\text{AgIoU}=\frac{2\cdot\text{F1}}{2-\text{F1}}.$
-Appendix E Evaluation and Comparison
-E.1 Evaluation
-Figure 7: Metrics across thresholds on the test set of 100 antigens.
-BConformeR was evaluated at various thresholds on 100 antigen records.
-As shown in Figure 7, the range between 0.1 and 0.5—highlighted by gray dashed lines—represents the region where precision, recall, F1 score and AgIoU perform well.
-Notably, the F1 score peaks around 0.44 when thresholds are set between 0.1 and 0.3.
-E.2 Baseline methods
-Thresholds for all baselines were adjusted to reflect their best scoring behavior, ensuring a fair comparison with BConformeR in the blind test.
-We observed that SEPPA-3.0 performs better under a low decision threshold (around 0.1).
-For SEMA-1D 2.0, which defines epitope residues using an 8 Å distance cutoff, the predicted scores tend to be higher; thus, we adopted a threshold of 0.5 above its default of 0.351.
-Moreover, although the default threshold of BepiPred-3.0 is 0.1512, we found that a slightly higher value around 0.22 led to better balance between precision and recall on our test set.
-Appendix F Ablation
-F.1 ViT1D
-Model Embed Dimension Seq Length Patch Size Depth Heads MLP Ratio
-ViT1D-9 1536 1024 1 9 12 2.0
-ViT1D-12 1536 1024 1 12 12 2.0
+Appendix C Ablation
+C.1 ViT1D
+Model Embed Dim Patch Size Depth Heads MLP Ratio
+ViT1D-9 1536 1 9 12 2.0
+ViT1D-12 1536 1 12 12 2.0
 Table 4: ViT variants used in ablation study.
@@
 The only difference is the number of transformer encoder layers.
-F.2 ResNet1D
+C.2 ResNet1D
 Stage Output Channels Kernel Size ResNet1D-101 ResNet1D-152
-input 1280 - $\checkmark$ $\checkmark$
-stem 320 7 $\checkmark$ $\checkmark$
-layer1 640 (160 $\times$ 4) 3 3 3
-layer2 1280 (320 $\times$ 4) 3 4 8
-layer3 2560 (640 $\times$ 4) 3 23 36
-layer4 5120 (1280 $\times$ 4) 3 3 3
+input 1152 - $\checkmark$ $\checkmark$
+stem 288 7 $\checkmark$ $\checkmark$
+layer1 576 (144 $\times$ 4) 3 3 3
+layer2 1152 (288 $\times$ 4) 3 4 8
+layer3 2304 (576 $\times$ 4) 3 23 36
+layer4 4608 (1152 $\times$ 4) 3 3 3
 classifier 2 1 $\checkmark$ $\checkmark$
@@
 We employed two variants of the ResNet1D backbone in the ablation studies: ResNet1D-101 and ResNet1D-152.
-As shown in Table 5, both models accept input feature maps with 1280 channels and begin with an initial convolutional layer producing 320 channels.
+As shown in Table 5, both models accept input feature maps with 1152 channels and begin with an initial convolutional layer producing 320 channels.
 The networks consist of four residual stages of bottlenecks with the expansion factor of 4, each stage starting by downsampling the channel dimension to half.
@@
 The final output is projected to the target classes through a $\times1$ convolutional layer.
+.